### v1

# Patterns of Non-Stationary Solutions to Symmetric Systems of Second-Order Differential Equations

Abstract
We establish the existence of non-stationary solutions to a symmetric system of second-order autonomous differential equations. Our technique is based on the equivariant degree theory and involves a novel characterization of orbit types of maximal kind in the Burnside Ring product of a finite number of basic degrees for the group $O(2)\times\Gamma\times{\mathbb{Z}}_{2}$.
Mathematics Subject Classification: Primary: 34C25, 37C81, 47H11, 55M25, 19A22
Key Words and Phrases: symmetric equation, existence of solutions, equivariant Leray-Schauder degree, nonlinear analysis, wave equation.
1 Introduction
One of the paradigmatic problems used to validate novel techniques in nonlinear analysis is classical forced pendulum equation $\ddot{x}=\bm{f}(x)$ (cf. [7]-[11]). More generally, dynamical systems described by second-order differential equations strike a delicate balance between simplicity and a capacity for capturing rich and intricate nonlinear behavior, making them ideal candidates for the testing of analytical methods. On the other hand, autonomous differential equations admit natural $SO(2)$ symmetries related to time shifting that make them particularly well-suited for a framework based on the equivariant degree theory. Given a finite group $\Gamma$ and a fixed number $p>0$, our primary objective is to demonstrate how the Leray-Schauder equivariant degree can be used to establish the existence of non-stationary solutions to the $\Gamma$-symmetric network of wave equations
$\begin{cases}\ddot{x}(t)=f(x(t))+Ax(t),\quad t\in\mathbb{R},\;x(t)\in V\\ x(t)=x(t+p),\;\dot{x}(t)=\dot{x}(t+p),\end{cases}$ (1)
where $V:=\mathbb{R}^{N}$ is an orthogonal $\Gamma$-representation, $A:V\rightarrow V$ is a symmetric $\Gamma$-equivariant matrix and $f:V\rightarrow V$ is a continuous function satisfying the conditions
  1. ($A_{1}$)
    $f$ is $\Gamma$-equivariant, i.e. $f(\gamma x)=\gamma f(x)$ for all $\gamma\in\Gamma,\;t\in\mathbb{R},\;x\in V$.
  2. ($A_{2}$)
    $f$ is odd, i.e. $f(-x)=-f(x)$ for all $x\in V$.
  3. ($A_{3}$)
    the Jacobian of $f$ at the origin $Df(0):V\rightarrow V$ is trivial, i.e. for all $x\in V$ one has
    $Df(0)x=0.$
  4. ($A_{4}$)
    there exists a constant $M>0$ such that for all $x\in V$ with $|x|\geq M$, one has
    $f(x)\cdot x>0.$
In the context of an arrangement of $N$ identical pendula, subjected to nonlinear forcing represented by $f$ and with coupling relations described by the matrix $A$, the vectors
$\begin{cases}x(t)=(x_{1}(t),x_{2}(t),\ldots,x_{N}(t)),\\ \dot{x}(t)=(\dot{x}_{1}(t),\dot{x}_{2}(t),\ldots,\dot{x}_{N}(t));\\ \ddot{x}(t)=(\ddot{x}_{1}(t),\ddot{x}_{2}(t),\ldots,\ddot{x}_{N}(t)),\end{cases}$
contain the angles of the pendula from the vertical position, their velocities and their accelerations at any time $t>0$, respectively. The effectiveness of the equivariant Leray-Schauder degree in detecting solutions to a system of a similar form to (1) for several explicit choices of the symmetry group $\Gamma$ has already been demonstrated by Garcia-Azpetia et al. in the article [5]. Our contribution to this problem is a practical existence result that applies for any arbitrary choice of $\Gamma$ and our technique is broadly applicable to any model exhibiting symmetries of the product group
$G:=O(2)\times\Gamma\times{\mathbb{Z}}_{2}.$
1.1 Setting for the Leray-Schauder Equivariant Degree
With the substitutions $\beta:=p/2\pi$ and $u(t):=x(\beta t)$, the equations (1) become
$\begin{cases}\ddot{u}(t)=\beta^{2}f(u(t))+\beta^{2}Au(t),\quad t\in\mathbb{R},\;u(t)\in V\\ u(t)=u(t+2\pi),\;\dot{u}(t)=\dot{u}(t+2\pi).\end{cases}$ (2)
Notice that any $2\pi$-periodic solution $u(t):\mathbb{R}\rightarrow V$ for the system (2) corresponds to the $p$-periodic solution $x(t)=u(t/\beta)$ for our original system (1). We denote by $\mathscr{E}:=C_{2\pi}(\mathbb{R},V)$ the Banach space of all continuous, $2\pi$-periodic, $V$-valued functions with the standard norm
$\|\varphi\|_{\infty}:=\max_{t\in\mathbb{R}}|\varphi(t)|,\quad\varphi\in C_{2\pi}(\mathbb{R},V),$
and by $\mathscr{H}:=C^{2}_{2\pi}(\mathbb{R},V)$ the subspace of twice continuously differentiable functions equipped with the norm
$\|u\|_{\mathscr{H}}:=\max\{\|u\|_{\infty},\|\dot{u}\|_{\infty},\|\ddot{u}\|_{\infty}\},\quad u\in\mathscr{H}.$
Following [5], we also consider the differential operator
$\mathscr{L}:\mathscr{H}\rightarrow\mathscr{E},\quad\mathscr{L}u:=\ddot{u}-u,$
the Nemystskii operator
$N:\mathscr{E}\rightarrow\mathscr{E},\quad(Nu)(t):=f(u(t)),$
and the Banach embedding
$j:\mathscr{H}\rightarrow\mathscr{E},\quad j(u)(t):=u(t).$
Since $\mathscr{L}:\mathscr{H}\rightarrow\mathscr{E}$ is a linear isomorphism, $N:\mathscr{E}\rightarrow\mathscr{E}$ is continuous and $j:\mathscr{H}\rightarrow\mathscr{E}$ is compact, the operator $\mathscr{F}:\mathscr{H}\rightarrow\mathscr{H}$ given by
$\mathscr{F}(u):=u-\mathscr{L}^{-1}(\beta^{2}N(j(u))+\beta^{2}Aj(u)-j(u)),$ (3)
is a compact perturbation of the identity on $\mathscr{H}$ and the system (2) is equivalent to the operator equation
$\mathscr{F}(u)=0,$ (4)
in the sense that a function $u\in\mathscr{H}$ is a solution to (2) if and only if it satisfies (4). In what follows, (2) will be called the operator equation associated with the system (2). Notice that the space $\mathscr{H}$ is an isometric Banach $G$-representation with respect to the $G$-action $G\times\mathscr{H}\rightarrow\mathscr{H}$
$(e^{i\theta},\gamma,\pm 1)u(t):=\pm\gamma u(t+\theta),\quad(\kappa,\gamma,\pm 1)u(t):=\pm\gamma u(-t),\quad u\in\mathscr{H},\;t\geq 0.$
where $(e^{i\theta},\gamma,\pm 1),(\kappa,\gamma,\pm 1)\in O(2)\times\Gamma\times{\mathbb{Z}}_{2}$. Gracia-Azpetia et al. demonstrate the applicability of the $G$-equivariant Leray-Schauder degree to (4) in [5] by proving that the assumptions ($A_{1}$)($A_{4}$) imply that nonlinear operator (3) is a completely continuous $G$-equivariant field, differentiable at the origin $0\in\mathscr{H}$ with
$\mathscr{A}u=\operatorname{Id}-\mathscr{L}^{-1}(\beta^{2}Au-u),\quad\mathscr{A}:=D_{u}\mathscr{F}(0):\mathscr{H}\rightarrow\mathscr{H}.$
For proof of the following Lemma, the reader is referred to [2, 3].
Lemma 1.1.
Under the conditions ($A_{1}$)-($A_{4}$), the operator $\mathscr{F}:\mathscr{H}\rightarrow\mathscr{H}$ is $\Omega:=B_{R}(0)\setminus B_{\epsilon}(0)$-admissible for some sufficiently small choice of $\epsilon>0$ and a sufficiently large choice of $R>0$. Moreover, one has
$G\text{\rm-deg}(\mathscr{F},\Omega)=(G)-G\text{\rm-deg}(\mathscr{A},B(\mathscr{H})),$
where $(G)\in A(G)$ is the unit element of the Burnside Ring.
Lemma (1.1) provides a sufficient condition for the existence of a non-trivial solution to (2); namely, if one has
$G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))\neq(G),$ (5)
then there exists a function $u\in\mathscr{H}\setminus\{0\}$ satisfying (2). A more practical reformulation of the condition (5): if for some orbit type $(H)\in\Phi_{0}(G)\setminus\{(G)\}$ one has
$\operatorname{coeff}^{H}\left(G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))\right)\neq 0,$
then, by the existence property of the $G$-equivariant Leray-Schauder degree (cf. Appendix A), there exists a map $u\in\mathbb{R}$ with an isotropy subgroup $G_{u}\leq G$ satisfying $(G_{u})\geq(H)$.
1.2 Towards a Computational Formula for $G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))$
Without specifying the group $\Gamma$, we assume that a complete list of the irreducible $\Gamma$-representations $\{\mathcal{V}_{j}\}_{j=0}^{r}$ is made available. As a $\Gamma$-representation, the space $V$ is also a natural $\Gamma\times\mathbb{Z}_{2}$-representation with the $\Gamma\times\mathbb{Z}_{2}$-isotypic decomposition
$V=V_{0}\oplus V_{1}\oplus\cdots\oplus V_{r},$
where each $\Gamma\times\mathbb{Z}_{2}$-isotypic component $V_{j}$ is modeled on the irreducible $\Gamma\times\mathbb{Z}_{2}$-representation $\mathcal{V}_{j}^{-}$ (here the superscript is meant to indicate that $\mathcal{V}_{j}$ has been equipped with antipodal $\mathbb{Z}_{2}$-action) in the sense that $V_{j}$ is equivalent to the direct sum of some finite number of copies of $\mathcal{V}_{j}^{-}$, i.e.
$V_{j}\simeq\mathcal{V}_{j}^{-}\oplus\cdots\oplus\mathcal{V}_{j}^{-}.$
The exact number of irreducible $\Gamma$-representations $\mathcal{V}_{j}^{-}$ ‘contained’ in the $\Gamma\times\mathbb{Z}_{2}$-isotypic component $V_{j}$ is called the $\mathcal{V}_{j}$-isotypic multiplicity of $V$ and is calculated according to the ratio,
$m_{j}:=\dim V_{j}/\dim\mathcal{V}_{j}^{-},\quad j\in\{0,1,\ldots,r\}.$ (6)
Since $A:V\rightarrow V$ is $\Gamma$-equivariant, one has
$\sigma(A)=\bigcup_{j=0}^{r}\sigma(A_{j}),\quad A_{j}:=A|_{V_{j}}:V_{j}\rightarrow V_{j}.$
In order to simplify our analysis, we enforce an additional non-degeneracy assumption:
  1. ($A_{0}$)
    for each $\Gamma$-isotypic component $V_{j}$, there exists a number $\mu_{j}\in\mathbb{R}$ with $\mu_{j}\neq-m^{2}/\beta^{2}$ for all $m\in{\mathbb{N}}$ such that
    $A_{j}=\mu_{j}\operatorname{Id}|_{V_{j}}:V_{j}\rightarrow V_{j}.$
Remark 1.2.
In [5], Garcia-Azpetia et al. avoid the assumption ($A_{0}$) with an appropriate fixed point reduction. For the sake of simplicity, we choose to sidestep the fixed point argument by assuming ($A_{0}$).
For each $s\in{\mathbb{N}}$, we denote by $\mathcal{W}_{s}\simeq\mathbb{C}$ the irreducible $O(2)$-representation equipped with the $s$-folding $O(2)$-action
$e^{i\theta}w:=e^{is\theta}w,\quad\kappa w:=\bar{w},\quad\theta\in SO(2),\;\kappa\in O(2),\;w\in\mathcal{W}_{s},$
and by $\mathcal{W}_{0}\simeq\mathbb{R}$ the irreducible $O(2)$-representation on which $O(2)$ acts trivially. Notice that the space $\mathscr{H}$ admits the $O(2)$-isotypic decomposition
$\mathscr{H}:=\overline{\bigoplus\limits_{m=0}^{\infty}\mathscr{H}_{m}},\quad\mathscr{H}_{m}:=\{u\in\mathscr{H}:u(t)=\cos(mt)a+\sin(mt)b,\;a,b\in V\},$
where
$\mathscr{H}_{m}\simeq\mathcal{W}_{m}\otimes V,\quad m\in{\mathbb{N}}\cup\{0\}.$
In the standard way, the $G$-isotypic decomposition of $\mathscr{H}$ can now be expressed in terms of $G$-isotypic components $\mathscr{H}_{m,j}\subset\mathscr{H}$ modeled on the irreducible $G$-representations
$\mathcal{V}_{m,j}:=\mathcal{W}_{m}\otimes\mathcal{V}_{j}^{-},\quad m\in{\mathbb{N}}\cup\{0\},\;j\in\{0,1,\ldots,r\},$
as follows
$\mathscr{H}=\bigoplus_{j=0}^{r}\overline{\bigoplus\limits_{m=0}^{\infty}\mathscr{H}_{m,j}},\quad\mathscr{H}_{m,j}:=\{u\in\mathscr{H}:u(t)=\cos(mt)a+\sin(mt)b,\;a,b\in V_{j}\}.$
With the $G$-isotypic decomposition of $\mathscr{H}$ at hand, we can begin to collect spectral data related to the $G$-equivariant linear operator $\mathscr{A}:\mathscr{H}\rightarrow\mathscr{H}$. For example, by Schur’s Lemma, one has
$\mathscr{A}(\mathscr{H}_{m,j})\subset\mathscr{H}_{m,j},\quad m\in{\mathbb{N}}\cup\{0\},\;j\in\{0,1,\ldots,r\}.$
such that, adopting the notation
$\mathscr{A}_{m,j}:=\mathscr{A}|_{\mathscr{H}_{m,j}}:\mathscr{H}_{m,j}\rightarrow\mathscr{H}_{m,j},$
the spectrum of $\mathscr{A}$ admits the following decomposition
$\sigma(\mathscr{A})=\bigcup\limits_{m=0}^{\infty}\bigcup\limits_{j=0}^{r}\sigma(\mathscr{A}_{m,j}).$
Under assumption ($A_{0}$), the spectrum of each $\mathscr{A}_{m,j}$ consists of the nonzero eigenvalue
$\mu_{m,j}:=1+\frac{\beta^{2}\mu_{j}-1}{1+m^{2}}=\frac{m^{2}+\beta^{2}\mu_{j}}{1+m^{2}},$
with multiplicity $m_{j}\in{\mathbb{N}}$ (cf. discussion of isotypic multiplicity (6)). The product property of the Leray-Schauder $G$-equivariant degree (cf. Appendix A) permits us to express the degree $G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))$ in terms of a Burnside Ring product of the Leray-Schauder $G$-equivariant degrees of the various restrictions $\mathscr{A}_{m,j}:\mathscr{H}_{m,j}\rightarrow\mathscr{H}_{m,j}$ of the $G$-equivariant linear isomorphism $\mathscr{A}:\mathscr{H}\rightarrow\mathscr{H}$ to the $G$-isotypic components $\mathscr{H}_{m,j}$ on their respective open unit balls $B(\mathscr{H}_{m,j}):=\{u\in\mathscr{H}_{m,j}\;:\;\|u\|_{\mathscr{H}}<1\}$ as follows
$G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))=\prod\limits_{j=0}^{r}\prod\limits_{m=0}^{\infty}G\text{\rm-deg}(\mathscr{A}_{m,j},B(\mathscr{H}_{m,j})).$ (7)
On the other hand, each degree $G\text{\rm-deg}(\mathscr{A}_{m,j}(\alpha),B(\mathscr{H}_{m,j}))$ is fully specified by the $\mathcal{V}_{j}$-isotypic multiplicities $\{m_{j}\}_{j=0}^{r}$ together with the real spectra of $\mathscr{A}_{m,j}$ according to formula,
$G\text{\rm-deg}(\mathscr{A}_{m,j},B(\mathscr{H}_{m,j}))=\begin{cases}(\deg_{\mathcal{V}_{m,j}})^{m_{j}}\quad&\text{ if }\mu_{m,j}<0;\\ (G)\quad&\text{ otherwise,}\end{cases}$ (8)
where $\deg_{\mathcal{V}_{m,j}}\in A(G)$ is the basic degree associated with the irreducible $G$-representation $\mathcal{V}_{m,j}$ (cf. Appendix A) and $(G)\in A(G)$ is the unit element of the Burnside Ring. In addition, since each basic degree $\deg_{\mathcal{V}_{m,j}}$ is involutive in the Burnside Ring (cf. Appendix A), one has
$(\deg_{\mathcal{V}_{m,j}})^{m_{j}}=\begin{cases}\deg_{\mathcal{V}_{m,j}}\quad&\text{ if }2\nmid m_{j};\\ (G)\quad&\text{ otherwise.}\end{cases}$ (9)
Putting together (7), (8) and (9), we introduce some notations to keep track of the indices
$\Sigma:=\left\{(m,j):m\in\mathbb{N}\cup\{0\},\;j\in\{0,1,\ldots,r\}\right\},$
which which contribute non-trivially to $G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))$. To begin, the negative spectrum of $\mathscr{A}:\mathscr{H}\rightarrow\mathscr{H}$ is accounted for with the index set
$\Sigma_{-}:=\left\{(m,j)\in\Sigma:\mu_{m,j}<0\right\}.$
Combining this with formulas (7) and (8) yields
$G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))\;=\prod\limits_{(m,j)\in\Sigma_{-}}(\deg_{\mathcal{V}_{m,j}})^{m_{j}}.$
This computation can be further reduced by discarding the indices $(m,j)\in\Sigma$ associated with even $\mathcal{V}_{j}$-isotypic multiplicities $m_{j}$, whose corresponding basic degrees contribute trivially to the Burnside Product (7). Specifically, we put
[…]
which, together with (9), permits us to express $G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))$ as the Burnside Ring product of a finite number of unique basic degrees, as follows
$G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))\;=\prod\limits_{(m,j)\in\Sigma_{0}}\deg_{\mathcal{V}_{m,j}}.$ (10)
1.3 Orbit Types of Maximal Kind
It is always possible to study the non-triviality of the product of a finite number of Burnside Ring elements (such as the product of basic degrees in the expression (10)) as a problem concerning the non-triviality of the product of a finite number of generator pairs $(K),(L)\in\Phi_{0}(G)$. Keeping in mind that the non-triviality of
$(K)\cdot(L)\in A(G),$
is equivalent to the existence of an orbit type $(H)\in\Phi_{0}(G)\setminus\{(G)\}$ with $\operatorname{coeff}^{H}((K)\cdot(L))\neq 0$, notice that the recurrence formula for the Burnside Ring product (cf. Appendix A)
$\operatorname{coeff}^{H}((K)\cdot(L))=\frac{n(H,L)|W(L)|n(H,K)|W(K)|-\sum_{(\tilde{H})>(H)}\operatorname{coeff}^{\tilde{H}}((K)\cdot(L))n(H,\tilde{H})|W(\tilde{H})|}{|W(H)|},$
is only useful if one first identifies the set of all orbit types $(\tilde{H})\in\Phi_{0}(G)\setminus\{(G)\}$ with $(\tilde{H})>(H)$. Therefore, computation of the coefficients $\operatorname{coeff}^{H}((K)\cdot(L))$ is generally impractical except for those orbit types which are maximal in $\Phi_{0}(G)$, i.e. for those orbit types $(H)\in\Phi_{0}(G)$ where $(\tilde{H})>(H)$ implies $(\tilde{H})=(G)$, in which case the recurrence formula simplifies to
$\operatorname{coeff}^{H}((K)\cdot(L))=\frac{n(H,L)|W(L)|n(H,K)|W(K)|}{|W(H)|}.$
Remark 1.3.
One of the principal difficulties when working with a product group of the form $G=O(2)\times\Gamma\times{\mathbb{Z}}_{2}$ is that the conjugacy class of the subgroup $O(2)\times\{e_{\Gamma}\}\times\{1\}\leq G$ is always a generator of the Burnside Ring $A(G)$. Since any orbit type $(H)$ which is maximal in the isotropy lattice $\Phi_{0}(G;\mathscr{H}\setminus\{0\})$ must also be maximal in sublattice $\Phi_{0}(G;\mathscr{H}_{0}\setminus\{0\})$, this implies that any non-zero function $u\in\mathscr{H}$ with an isotropy group $G_{u}\leq G$ such that $(G_{u})$ is a maximal element in $\Phi_{0}(G;\mathscr{H}\setminus\{0\})$ must be $O(2)$-invariant. Therefore, any map $u\in\mathscr{H}$ which is both non-trivial and non-constant has an isotropy group $G_{u}\leq G$ with a conjugacy class that is maximal in $\Phi_{0}(G;\mathscr{H}_{m}\setminus\{0\})$ for some positive $m$.
Motivated by the Remark 1.3, the conjugacy class of a subgroup $H\leq G$ is called an orbit type of maximal kind if it is a maximal element in $\Phi_{0}(G;\mathscr{H}_{m}\setminus\{0\})$ for some $m>0$. Moreover, we denote by $\mathfrak{M}$ the set of all orbit types of maximal kind, by $\mathfrak{M}_{m}$ the set of maximal elements in $\Phi_{0}(G;\mathscr{H}_{m}\setminus\{0\})$ and by $\mathfrak{M}_{m,j}$ the set of orbit types $\mathfrak{M}_{m}\cap\Phi_{0}(G;\mathscr{H}_{m,j}\setminus\{0\})$. Since each orbit type of maximal kind $(H)\in\mathfrak{M}_{m}$ is contained in the sublattice of orbit types $\Phi_{0}(G;\mathscr{H}_{m,j}\setminus\{0\})$ for at least one $j\in\{0,1,\ldots,r\}$, we have
$\mathfrak{M}=\bigcup_{m>0}\mathfrak{M}_{m},\quad\mathfrak{M}_{m}=\bigcup_{j=0}^{r}\mathfrak{M}_{m,j},$
corresponding to the $\Gamma\times\mathbb{Z}_{2}$-isotypic decomposition of the $m$-th $O(2)$-isotypic component of $\mathscr{H}$:
$\mathscr{H}_{m}=\mathscr{H}_{m,1}\oplus\mathscr{H}_{m,2}\oplus\cdots\oplus\mathscr{H}_{m,r}.$
At this point, we can strengthen the sufficient condition for the existence of a non-trivial solution to (2) presented in Section 1.1 to a sufficient condition for the existence of a non-stationary solution; namely, if for some orbit type of maximal kind $(H)\in\mathfrak{M}$ one has
$\operatorname{coeff}^{H}\left(G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))\right)\neq 0,$
then there exists a non-stationary function $u\in\mathscr{H}$ satisfying (2).
1.4 The $s$-Folding Homomorphism
In this section, we introduce the concept of $s$-folding, which plays a key role in our study of orbit types of maximal kind. For each $s\in\mathbb{N}$, we define the $s$-folding homomorphism by
$\phi_{s}:O(2)\rightarrow O(2)/\mathbb{Z}_{s}\simeq O(2),\quad\phi_{s}(e^{i\theta}):=e^{is\theta},\quad\phi_{s}(\kappa e^{i\theta}):=\kappa e^{is\theta}.$
There is a natural extension of $\phi_{s}$ to the Lie group homomorphism $\psi_{s}:O(2)\times\Gamma\times\mathbb{Z}_{2}\rightarrow O(2)\times\Gamma\times\mathbb{Z}_{2}$ given by
$\psi_{s}(e^{i\theta},\gamma,\pm 1):=(\phi_{s}(e^{i\theta}),\gamma,\pm 1),\quad\psi_{s}(\kappa e^{i\theta},\gamma,\pm 1):=(\phi_{s}(\kappa e^{i\theta}),\gamma,\pm 1).$ (11)
In turn, (11) induces the Burnside Ring homomorphism $\Psi_{s}:A(G)\rightarrow A(G)$ defined as follows
$\Psi_{s}(H):=({}^{s}H),\quad{}^{s}H:=\psi^{-1}_{s}(H).$ (12)
Clearly, one has the following relation between the sets $\mathfrak{M}_{m}$ and $\mathfrak{M}_{sm}$ for all $s>0$ and $m\geq 0$
$\Psi_{s}(\mathfrak{M}_{m})=\mathfrak{M}_{sm}.$
Zalman et al. hypothesize in [4]—where it serves as an essential component in the proof of their main result—that every orbit type $(H)\in\Phi_{0}(G;\mathscr{H}_{m})\setminus\{(G)\}$ satisfies the relation
$({}^{s}H)>(H)\text{ for all }s\in{\mathbb{N}}.$ (13)
In this paper, we demonstrate that the relation (13) does not hold universally. Instead, we prove that the relation between $(H)$ and $({}^{s}H)$ depends both on $(H)\in\Phi_{0}(G;\mathscr{H}_{m})$ and $s\in{\mathbb{N}}$.
1.5 An Outline of the Paper
The remainder of this paper is organized as follows: In Section 2, we present a classification for the generators of the Burnside Ring $A(G)$ according to their lattice relations with respect to the $s$-folding homomorphism (12). This classification is then used, in Section 3, to derive a practical formula for characterizing the non-triviality of an orbit type of maximal kind in the product of a finite number of basic degrees. Specifically, we find that the contribution of each eigenvalue $\mu_{m,j}\in\sigma(\mathscr{A})$ to the appearance of an orbit type of maximal kind $(H)\in\mathfrak{M}$ in the degree $G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))$ depends on $(i)$ the sign of $\mu_{m,j}$ and $(ii)$ the parity of the dimension of the corresponding $H$-fixed-point space $\mathcal{V}_{m,j}^{H}$. To formalize these findings, we introduce the following definition:
Definition 1.4.
Take $m>0$ and let $(H)$ be a maximal element in the isotropy lattice $\Phi_{0}(G;\mathscr{H}_{m}\setminus\{0\})$. An eigenvalue $\mu_{j}\in\sigma(A)$ is called $H$-essential if the following two relations are satisfied
$\mu_{j}<\frac{-m^{2}}{\beta^{2}},\quad 2\nmid\dim\mathcal{V}_{m,j}^{H}.$
Remark 1.5.
In the field of equivariant bifurcation theory, the Ize Conjecture (IC) refers to the proposition that every irreducible representation of a compact Lie group admits at least one subgroup with an odd dimensional fixed point space. Although the IC was shown to be false by Lauterbach et al. in [12] with counter examples of irreducible representations which only admit even dimensional fixed point spaces, the conjecture holds in many cases.
Finally, in Section 4, we demonstrate how these results lead to our main existence theorem for the system (1):
Theorem 1.6.
Take $m>0$ and let $(H)$ be a maximal element in the isotropy lattice $\Phi_{0}(G;\mathscr{H}_{m}\setminus\{0\})$. If the matrix $A:V\rightarrow V$ admits an odd number of $H$-essential eigenvalues, then there exists a non-stationary solution $u\in\mathscr{H}\setminus\{0\}$ to the system (1) with an isotropy subgroup $G_{u}\leq G$ satisfying $(G_{u})\geq(H)$.
2 A Classification of Generators for the Burnside Ring Associated with the Group $G=O(2)\times\Gamma\times\mathbb{Z}_{2}$.
In order to proceed with our computations in the Burnside Ring, we must employ the convention of amalgamated notation: a shorthand for the specification of subgroups in a product group, first considered by Balanov et al. in [3]. It is a well known consequence of Goursat’s Lemma (cf. [6]) that any closed subgroup in the product group $G=O(2)\times\Gamma\times{\mathbb{Z}}_{2}$ can be identified, up to its conjugacy class in $\Phi_{0}(G)$, with a quintuple $(K_{O},K_{\Gamma},L,\varphi_{O},\varphi_{\Gamma})$ consisting of two subgroups $K_{O}\leq O(2)$, $K_{\Gamma}\leq\Gamma\times\mathbb{Z}_{2}$, a group $L$ and a pair of epimorphisms $\varphi_{O}:K_{O}\rightarrow L$, $\varphi_{\Gamma}:K_{\Gamma}\rightarrow L$, as follows
$K_{O}{}^{\varphi_{O}}\times^{\varphi_{\Gamma}}_{L}K_{\Gamma}:=\{(x,y)\in K_{O}\times K_{\Gamma}\;:\;\varphi_{O}(x)=\varphi_{\Gamma}(y)\}.$ (14)
Let $H$ be a closed subgroup in $G$ with the amalgamated decomposition (14) and denote by $r(H)\in SO(2)$ the rotation generator in $K_{O}$. Since the order of $\varphi_{O}(r(H))\in L$ is invariant under conjugation of $r(H)$ in $K_{O}$, the following map is well-defined:
$\mathfrak{m}:\Phi_{0}(G)\rightarrow{\mathbb{N}},\quad\mathfrak{m}(H):=|\varphi_{O}(r(H))|.$ (15)
Given any orbit type $(H)\in\Phi_{0}(G)$ and a number $s_{0}\in{\mathbb{N}}$, we will demonstrate how the map (15) can be used to to describe the set
$\mathfrak{N}(H,s_{0}):=\{s\in{\mathbb{N}}:({}^{s}H)\leq({}^{s_{0}}H)\}.$ (16)
In turn, for any orbit type of maximal kind $(H)\in\mathfrak{M}$ and for any pair of generators $(K),(L)\in\Phi_{0}(G)$ we can use the set (16) to characterize the coefficients of $({}^{s}H)$ for all $s\in{\mathbb{N}}$ in the Burnside Ring product
$(K)\cdot(L)\in A(G).$
Remark 2.1.
For any orbit type of maximal kind $(H)\in\mathfrak{M}$, the relation $(H)<(\tilde{H})$ implies $(\tilde{H})=(G)$ or $(\tilde{H})=({}^{s}H)$ for some $s\in{\mathbb{N}}$. Therefore, since satisfaction of the relations $(H)<(K)$ and $(H)<(L)$ is a prerequisite for the non-triviality $\operatorname{coeff}^{H}((K)\cdot(L))\neq 0$, we can restrict our focus to the coefficients of $({}^{s}H)$ for all $s\in{\mathbb{N}}$ in Burnside Ring products of the form
$({}^{s_{0}}H)\cdot({}^{s_{1}}H)\in A(G),\quad s_{0},s_{1}\in{\mathbb{N}}.$
Theorem 2.2.
For any orbit type $(H)\in\Phi_{0}(G)$ and pair of numbers $s_{0},s_{1}\in{\mathbb{N}}$ with $s_{1}\leq s_{0}$, the corresponding pair of folded orbit types $({}^{s_{1}}H),({}^{s_{0}}H)\in\Phi_{0}(G)$ have the relation $({}^{s_{1}}H)\leq({}^{s_{0}}H)$ if and only if
$\mathfrak{m}(H)\mid s_{0}-s_{1}\text{ or }\mathfrak{m}(H)\mid s_{0}+s_{1}.$
Proof.
Assume that a subgroup representative $H\in(H)$ has the amalgamated decomposition $K_{O}{}^{\varphi_{O}}\times^{\varphi_{\Gamma}}_{L}K_{\Gamma}$. According to the natural ordering of $\Phi_{0}(G)$, the relation
$({}^{s_{1}}H)\leq({}^{s_{0}}H)$ (17)
is equivalent to the existence of an element $g\in G$ for which
$g{}^{s_{1}}Hg^{-1}\leq{}^{s_{0}}H.$ (18)
Given any $s\in{\mathbb{N}}$, put ${}^{s}K_{O}:=\phi_{s}^{-1}(K_{O})$, $\varphi^{s}_{O}:=\varphi_{O}\circ\phi_{s}:{}^{s}K_{O}\rightarrow L$ and notice that the subgroup ${}^{s}H\leq G$ has the amalgamated decomposition
$\psi_{s}^{-1}(K_{O}{}^{\varphi_{O}}\times^{\varphi_{\Gamma}}_{L}K_{\Gamma})$ $=\psi_{s}^{-1}(\{(x,y)\in K_{O}\times K_{\Gamma}\;:\;\varphi_{O}(x)=\varphi_{\Gamma}(y)\})$
=$=\{(\phi_{s}^{-1}(x),y)\in\phi_{s}^{-1}(K_{O})\times K_{\Gamma}\;:\;\varphi_{O}(x)=\varphi_{\Gamma}(y)\}$
=$=\{(x^{\prime},y)\in{}^{s}K_{O}\times K_{\Gamma}\;:\;\varphi^{s}_{O}(x^{\prime})=\varphi_{\Gamma}(y)\}$
$={}^{s}K_{O}{}^{\varphi^{s}_{O}}\times^{\varphi_{\Gamma}}_{L}K_{\Gamma}.$
On the other hand, and without loss of generality, take $g:=(a,e_{\Gamma},1)\in O(2)\times\Gamma\times\mathbb{Z}_{2}$, put $\tilde{K}_{O}:=aK_{O}a^{-1}\leq O(2)$ and define the epimorphism $\tilde{\varphi}_{O}:\tilde{K}_{O}\rightarrow L$ by
$\tilde{\varphi}_{O}(x):=\varphi_{O}(a^{-1}xa),\quad x\in\tilde{K}_{O},$
so that the subgroup $gHg^{-1}\leq G$ can be expressed, using amalgamated notation, as follows
$gK_{O}{}^{\varphi_{O}}\times^{\varphi_{\Gamma}}_{L}K_{\Gamma}g^{-1}$ $=(a,e_{\Gamma},1)\{(x,y)\in K_{O}\times K_{\Gamma}\;:\;\varphi_{O}(x)=\varphi_{\Gamma}(y)\}(a^{-1},e_{\Gamma},1)$
=$=\{(axa^{-1},y)\in aK_{O}a^{-1}\times K_{\Gamma}\;:\;\varphi_{O}(x)=\varphi_{\Gamma}(y)\}$
=$=\{(\tilde{x},y)\in\tilde{K}_{O}\times K_{\Gamma}\;:\;\tilde{\varphi}_{O}(\tilde{x})=\varphi_{\Gamma}(y)\}$
$=\tilde{K}_{O}{}^{\tilde{\varphi}_{O}}\times^{\varphi_{\Gamma}}_{L}K_{\Gamma}.$
With these notations, it becomes clear that the subgroups ${}^{s_{0}}H,{}^{s_{1}}H\leq G$ have the relation (18) for some $g=(a,\gamma,\pm 1)\in O(2)\times\Gamma\times{\mathbb{Z}}_{2}$ if and only if $\varphi^{s_{0}}_{O}(axa^{-1})=\varphi^{s_{1}}_{O}(x)$ for every $(x,y)\in K_{O}\times K_{\Gamma}$ with $\varphi_{O}(x)=\varphi_{\Gamma}(y)$ such that the relation (17) is equivalent to the existence of an element $a\in O(2)$ satisfying
$\varphi^{s_{1}}_{O}(axa^{-1})=\varphi^{s_{0}}_{O}(x)\text{ for all }x\in K_{O}.$
Clearly, it is sufficient to consider the behaviour of the rotation generator in $r\in K_{0}$ and the two cases $a\in\{\kappa,e_{O(2)}\}$. If $a=\kappa$, then from $\kappa r\kappa=r^{-1}$, one has
$\kappa{}^{s_{1}}H\kappa\leq{}^{s_{0}}H\iff\varphi_{O}^{s_{1}}(\kappa r\kappa)=\varphi_{O}^{s_{0}}(r)\iff\varphi_{O}(r)^{s_{0}+s_{1}}=e_{L}\iff\mathfrak{m}(H)\mid s_{0}+s_{1}.$
Supposing instead that $a=e_{O(2)}$, one has
${}^{s_{1}}H\leq{}^{s_{0}}H\iff\varphi_{O}^{s_{1}}(r)=\varphi_{O}^{s_{0}}(r)\iff\varphi_{O}(r)^{s_{0}-s_{1}}=e_{L}\iff\mathfrak{m}(H)\mid s_{0}-s_{1}.$
$\square$ For the sake of simplifying our exposition throughout the remainder of this paper, we employ the Iverson brackets which map any logical predicate $P$ to the set $\{0,1\}$ according to the rule
$[P]:=\begin{cases}1\quad&\text{ if }P\text{ is true};\\ 0\quad&\text{ otherwise}.\end{cases}$ (19)
Theorem 2.3.
Let $(H)\in\mathfrak{M}$ be an orbit type of maximal kind. For any pair of numbers $s_{0},s_{1}\in{\mathbb{N}}$, one has
$\operatorname{coeff}^{{}^{s}H}(({}^{s_{0}}H)\cdot({}^{s_{1}}H))=\begin{cases}|W(H)|\quad&\text{ if }s=\gcd(s_{0},s_{1})\text{ and }s\in\mathfrak{N}(H,s_{0})\cap\mathfrak{N}(H,s_{1});\\ 0\quad&\text{otherwise},\end{cases}$
equivalently, for any $s\in{\mathbb{N}}$, one has
$\operatorname{coeff}^{{}^{s}H}(({}^{s_{0}}H)\cdot({}^{s_{1}}H))=|W(H)|[s=\gcd{(s_{0},s_{1})}][s\in\mathfrak{N}(H,s_{0})\cap\mathfrak{N}(H,s_{1})].$
Proof.
Recall, the integer coefficient $n_{L}\in{\mathbb{Z}}$ of an orbit type $(L)$ in the Burnside Ring product $({}^{s_{0}}H)\cdot({}^{s_{1}}H)$ is defined as the number of orbits of type $(L)$ contained in the $G$-space $G/{}^{s_{0}}H\times G/{}^{s_{1}}H$. In particular, $(L)\in\Phi_{0}(G)$ is an orbit type in $G/{}^{s_{0}}H\times G/{}^{s_{1}}H$ if and only if there exists a group element $g\in G$ such that
$L={}^{s_{0}}H\cap g{}^{s_{1}}Hg^{-1}.$
Without loss of generality, put $g=(a,e_{\Gamma},1)$ with an arbitrary $a\in O(2)$ and notice that the amalgamated decomposition
${}^{s_{0}}H\cap g{}^{s_{1}}Hg^{-1}$ =$=\left\{(x,y)\in\left({}^{s_{0}}K_{O}\cap a{}^{s_{1}}K_{O}a^{-1}\right)\times K_{\Gamma}:\varphi_{O}^{s_{0}}(x)=\varphi_{O}^{s_{1}}(a^{-1}xa)=\varphi_{\Gamma}(y)\right\}$
$=\left({}^{s_{0}}K_{O}\cap{}^{s_{1}}\tilde{K}_{O}\right){}^{\psi}\times_{L}^{\varphi_{\Gamma}}K_{\Gamma},\quad(\text{where }{}^{s_{1}}\tilde{K}_{O}:=a{}^{s_{1}}K_{O}a^{-1})$
describes a subgroup belonging to the conjugacy class $({}^{s}H)$ if and only if there exists a group element $b\in O(2)$ with both
  1. (i)
    $\varphi_{O}(bxb^{-1})=\varphi_{O}^{s_{0}}(x)=\varphi_{O}^{s_{1}}(axa^{-1});\text{ for all }x\in K_{O},$
  2. (ii)
    $b{}^{s}K_{O}b^{-1}={}^{s_{0}}K_{O}\cap a{}^{s_{1}}K_{O}a^{-1}.$
The equivalence of condition $(\rm i)$ with the inclusion $s\in\mathfrak{N}(H,s_{0})\cap\mathfrak{N}(H,s_{1})$ has already been demonstrated in the proof of Theorem 2.2. On the other hand, the observation
$K_{O}\in\{O(2),SO(2),D_{n},{\mathbb{Z}}_{n},n\in{\mathbb{N}}\},$
together with the facts
${\mathbb{Z}}_{s_{0}n}\cap{\mathbb{Z}}_{s_{1}n}={\mathbb{Z}}_{\gcd(s_{1},s_{0})n}\text{ and }D_{n}={\mathbb{Z}}_{n}\cup\kappa{\mathbb{Z}}_{n},$
implies that condition $(\rm ii)$ can be equated with $s=\gcd(s_{0},s_{1})$ (cf. [2], Example 3.12). For convenience, put $d:=\gcd(s_{0},s_{1})$ and notice that, in the case one has $d\in\mathfrak{N}(H,s_{0})\cap\mathfrak{N}(H,s_{1})$, the coefficient $n_{d}\in{\mathbb{Z}}$ standing next to the orbit type $({}^{d}H)$ in the Burnside Ring product $({}^{s_{0}}H)\cdot({}^{s_{1}}H)\in A(G)$ can be obtained with the recurrence formula (cf. Appendix A) as follows
$n_{d}=\frac{n({}^{d}H,{}^{s_{0}}H)|W({}^{s_{0}}H)|n({}^{d}H,{}^{s_{1}}H)|W({}^{s_{1}}H)|}{|W({}^{d}H)|},$
and the result follows from the fact $|W(H)|=|W({}^{s}H)|$ for any $s\in{\mathbb{N}}$ together with the observation that $d\in\mathfrak{N}(H,s_{0})\cap\mathfrak{N}(H,s_{1})$ implies $n({}^{d}H,{}^{s_{0}}H)=n({}^{d}H,{}^{s_{1}}H)=1$. $\square$
3 A Characterization of the Orbit Types of Maximal Kind in the Product of Basic Degrees.
Since, the classification put forward in the previous section fully characterizes the orbit types $(\tilde{H})$ with $(\tilde{H})>(H)$ for any orbit type of maximal kind $(H)\in\mathfrak{M}$, we are now in a position to calculate their coefficients in Burnside Ring products of the form (10). Given a basic degree $\deg_{\mathcal{V}_{m,j}}\in A(G)$ corresponding to an index pair $(m,j)\in\Sigma$ with $m>0$ and an orbit type of maximal kind $(H)\in\mathfrak{M}$, the recurrence formula for the Leray-Schauder $G$-equivariant degree (cf. Appendix A) implies
$\deg_{\mathcal{V}_{m,j}}=(G)-y_{m,j}(H)+a_{m,j},$ (20)
where $a_{m,j}\in A(G)$ is such that $\operatorname{coeff}^{{}^{s}H}(a_{m,j})=0$ for all $s\in\mathbb{N}$ and the coefficient $y_{m,j}\in\mathbb{Z}$ is determined by the rule
$y_{m,j}=\begin{cases}0\quad&\text{if }\dim{\mathcal{V}_{m,j}^{H}}\text{ is even};\\ 1\quad&\text{if }\dim{\mathcal{V}_{m,j}^{H}}\text{ is odd and }|W(H)|=2;\\ 2\quad&\text{if }\dim{\mathcal{V}_{m,j}^{H}}\text{ is odd and }|W(H)|=1,\end{cases}$ (21)
or, equivalently (see discussion of the Iverson bracket notation (19)):
$y_{m,j}=x_{0}\left[2\nmid\dim{\mathcal{V}_{m,j}^{H}}\right],$
where
$x_{0}=\begin{cases}1\quad&\text{if }|W(H)|=2;\\ 2\quad&\text{if }|W(H)|=1.\end{cases}$
Lemma 3.1.
Let $(H)\in\mathfrak{M}$ be an orbit type of maximal kind. For any basic degree $\deg_{\mathcal{V}_{m,j}}$ with $(m,j)\in\Sigma$ and $m>0$ one has the implication
$\operatorname{coeff}^{H}(\deg_{\mathcal{V}_{m,j}})\neq 0\implies(H)\in\mathfrak{M}_{m,j}.$
Proof.
Since $\mathcal{V}_{m,j}^{H}=\{0\}$ for any orbit type $(H)\notin\Phi_{0}(G,\mathscr{H}_{m,j})$, the proof follows from formula (21) together with the assumption that $(H)$ is of maximal kind. $\square$ For every $s\in\mathbb{N}$, the $s$-folding homomorphism (11) induces the following relation between the basic degrees $\deg_{\mathcal{V}_{m,j}},\deg_{\mathcal{V}_{sm,j}}\in A(G)$
$\Psi_{s}(\deg_{\mathcal{V}_{m,j}})=\deg_{\mathcal{V}_{sm,j}},$ (22)
such that, for any orbit type $(H)\in\Phi_{0}(G)$, one has
$\operatorname{coeff}^{H}(\deg_{\mathcal{V}_{m,j}})=\operatorname{coeff}^{{}^{s}H}(\deg_{\mathcal{V}_{sm,j}}).$ (23)
Using the observations (20), (22), and (23), we will derive results which characterize the behavior of orbit types of maximal kind in the Burnside Ring product of a finite collection of basic degrees of the form
$\left\{\deg_{\mathcal{V}_{s_{k}m,j_{k}}}\in A(G):s_{k}\in\mathbb{N}\right\}_{k=1,2,\ldots,N},\quad m>0,\;j_{k}\in\{0,1,\ldots,r\}.$ (24)
To begin with, we recall a result that the coefficient of an orbit type of maximal kind $(H)\in\mathfrak{M}$ is $2$-nilpotent with respect to the Burnside Ring product of any pair of basic degrees $\deg_{\mathcal{V}_{m,j_{1}}},\deg_{\mathcal{V}_{m,j_{2}}}\in A(G)$ with $m>0$ and $j_{1},j_{2}\in\{0,1,\ldots,r\}$, provided that both $\dim\mathcal{V}_{m,j_{1}}^{H}$ and $\dim\mathcal{V}_{m,j_{2}}^{H}$ are odd. Although this fact has been demonstrated elsewhere (cf. for example [2], [3], [4], [5]), the method of its proof is instructive for the proofs of subsequent, novel results and for this reason we include our own version of the argument.
Lemma 3.2.
Let $(H)\in\mathfrak{M}$ be an orbit type of maximal kind. For any $m>0$ and $j_{1},j_{2}\in\{0,1,\ldots,r\}$, one has
$\operatorname{coeff}^{H}(\deg_{\mathcal{V}_{m,j_{1}}}\cdot\deg_{\mathcal{V}_{m,j_{2}}})=\begin{cases}0\quad&\text{if }\dim\mathcal{V}_{m,j_{1}}^{H}\text{ and }\dim\mathcal{V}_{m,j_{2}}^{H}\text{ are of the same parity;}\\ -x_{0}\quad&\text{else,}\end{cases}$
or, equivalently
$\operatorname{coeff}^{H}(\deg_{\mathcal{V}_{m,j_{1}}}\cdot\deg_{\mathcal{V}_{m,j_{2}}})=-x_{0}\left[2\nmid\dim\mathcal{V}_{m,j_{1}}^{H}+\dim\mathcal{V}_{m,j_{2}}^{H}\right].$
Proof.
Consider the Burnside Ring product of the relevant basic degrees
$\deg_{\mathcal{V}_{m,j_{1}}}\cdot\deg_{\mathcal{V}_{m,j_{2}}}$ $=\left((G)-y_{m,j_{1}}(H)+a_{m,j_{1}}\right)\cdot\left((G)-y_{m,j_{2}}(H)+a_{m,j_{2}}\right)$
$=(G)-(y_{m,j_{1}}+y_{m,j_{2}}-y_{m,j_{1}}y_{m,j_{2}}|W(H)|)(H)+a,$
where $a\in A(G)$ is such that $\operatorname{coeff}^{{}^{s}H}(a)=0$ for all $s\in\mathbb{N}$. Now, if $\dim\mathcal{V}_{m,j_{1}}^{H}$ and $\dim\mathcal{V}_{m,j_{2}}^{H}$ are both even, then one has $y_{m,j_{1}}=y_{m,j_{2}}=0$ and the result follows. On the other hand, if $\dim\mathcal{V}_{m,j_{1}}^{H}$ and $\dim\mathcal{V}_{m,j_{2}}^{H}$ are both odd, then one has $y_{m,j_{1}}=y_{m,j_{2}}=x_{0}$ such that
$y_{m,j_{1}}+y_{m,j_{2}}-y_{m,j_{1}}y_{m,j_{2}}|W(H)|=x_{0}(2-x_{0}|W(H)|),$
where in either of the cases: $x_{0}=2$ and $|W(H)|=1$ or $x_{0}=1$ and $|W(H)|=2$, one has $02-x_{0}|W(H)|=$ 0. Suppose instead that $\dim\mathcal{V}_{m,j_{1}}^{H}$ and $\dim\mathcal{V}_{m,j_{2}}^{H}$ are of different parties and one has the two corresponding cases, $y_{m,j_{1}}=x_{0}$ and $y_{m,j_{2}}=0$ or $y_{m,j_{1}}=0$ and $y_{m,j_{2}}=x_{0}$, both of which imply $y_{m,j_{1}}+y_{m,j_{2}}-y_{m,j_{1}}y_{m,j_{2}}|W(H)|=x_{0}$. $\square$ Since, our computations in the Burnside Ring often involve the product of a number of basic degrees, we consider a natural generalization of Lemma (3.2) to an orbit type of maximal kind $(H)\in\mathfrak{M}$ and the Burnside Ring product of any finite collection of basic degrees $\{\deg_{\mathcal{V}_{m,j_{k}}}\}_{k=1}^{N}$ with $m>0$ and $j_{1},\ldots,j_{N}\in\{0,1,\ldots,r\}$.
Corollary 3.3.
Let $(H)\in\mathfrak{M}$ be an orbit type of maximal kind. For any finite collection of basic degrees $\{\deg_{\mathcal{V}_{m,j_{k}}}\}_{k=1}^{N}$ with $m>0$ and $j_{1},\ldots,j_{N}\in\{0,1,\ldots,r\}$, one has
$\operatorname{coeff}^{H}\left(\prod\limits_{k=1}^{N}\deg_{\mathcal{V}_{m,j_{k}}}\right)=\begin{cases}0\quad&\text{if }\dim\mathcal{V}_{m,j_{k}}^{H}\text{ is odd for an even number of }j_{k}\in\{0,1,\ldots,r\};\\ -x_{0}\quad&\text{otherwise,}\end{cases}$
or, equivalently
$\operatorname{coeff}^{H}\left(\prod\limits_{k=1}^{N}\deg_{\mathcal{V}_{m,j_{k}}}\right)=-x_{0}\left[2\nmid\sum\limits_{k=1}^{N}\dim\mathcal{V}_{m,j_{k}}^{H}\right].$
The previous result only concerns the product of basic degrees associated with a fixed Fourier mode. We are also interested in characterizing the behaviour of the coefficients of an orbit type of maximal kind $(H)\in\mathfrak{M}$ and its $s$-foldings $\{({}^{s}H)\}_{s\in{\mathbb{N}}}\subset\mathfrak{M}$ in the Burnside Ring product of any pair of basic degrees $\deg_{\mathcal{V}_{s_{1}m,j_{1}}},\deg_{\mathcal{V}_{s_{2}m,j_{2}}}\in A(G)$ with $m>0$, $j_{1},j_{2}\in\{0,1,\ldots,r\}$ and where $s_{1},s_{2}\in{\mathbb{N}}$ are any two numbers satisfying $s_{1}\neq s_{2}$.
Lemma 3.4.
Let $(H)\in\mathfrak{M}$ be an orbit type of maximal kind. For any pair of basic degrees $\deg_{\mathcal{V}_{s_{1}m,j_{1}}},\deg_{\mathcal{V}_{s_{2}m,j_{2}}}\in A(G)$ with $m>0$, $j_{1},j_{2}\in\{0,1,\ldots,r\}$ and $s_{1},s_{2}\in{\mathbb{N}}$ where $s_{1}\neq s_{2}$ and for any number any $s_{0}\in{\mathbb{N}}$, one has
$\operatorname{coeff}^{{}^{s_{0}}H}$ $(\deg_{\mathcal{V}_{s_{1}m,j_{1}}}\cdot\deg_{\mathcal{V}_{s_{2}m,j_{2}}})=-x_{0}\left[2\nmid\dim\mathcal{V}_{m,j_{0}}^{H}\right]\left[s_{0}\in\{s_{1},s_{2}\}\right]$
$\quad+2x_{0}\left[2\nmid\dim\mathcal{V}_{m,j_{1}}^{H}\dim\mathcal{V}_{m,j_{2}}^{H}\right]\left[\operatorname{gcd}(s_{1},s_{2})=s_{0}\right][s_{0}\in\mathfrak{N}_{s_{1}}(H)\cap\mathfrak{N}_{s_{2}}(H)],$
where
$j_{0}:=\begin{cases}j_{1}&\text{ if }s_{0}=s_{1};\\ j_{2}&\text{ if }s_{0}=s_{2}.\end{cases}$
Proof.
Consider the Burnside Ring product of the relevant basic degrees
$\deg_{\mathcal{V}_{s_{1}m,j_{1}}}\cdot\deg_{\mathcal{V}_{s_{2}m,j_{2}}}$ $=\left((G)-y_{m,j_{1}}({}^{s_{1}}H)+a_{m,j_{1}}\right)\cdot\left((G)-y_{m,j_{2}}({}^{s_{2}}H)+a_{m,j_{2}}\right)$
$=(G)-y_{m,j_{1}}({}^{s_{1}}H)-y_{m,j_{2}}({}^{s_{2}}H)+y_{m,j_{1}}y_{m,j_{2}}({}^{s_{1}}H)\cdot({}^{s_{2}}H)+a$
$=(G)-y_{m,j_{1}}({}^{s_{1}}H)-y_{m,j_{2}}({}^{s_{2}}H)$
$\quad+y_{m,j_{1}}y_{m,j_{2}}|W(H)|({}^{\operatorname{gcd}(s_{1},s_{2})}H)[\operatorname{gcd}(s_{1},s_{2})\in\mathfrak{N}_{s_{1}}(H)\cap\mathfrak{N}_{s_{2}}(H)]+a$ (25)
where $a\in A(G)$ satisfies $\operatorname{coeff}^{{}^{s}H}(a)=0$ for all $s\in\mathbb{N}$. In order to demonstrate each of the boolean conditions, we consider the following cases:
  1. $(i)$
    Supposing that $s_{0}\notin\{s_{1},s_{2}\}$, if it is also the case that both $\dim\mathcal{V}_{m,j_{1}}^{H}$ and $\dim\mathcal{V}_{m,j_{2}}^{H}$ are odd and $\operatorname{gcd}(s_{1},s_{2})=s_{0}$, and $s_{0}\in\mathfrak{N}_{s_{1}}(H)\cap\mathfrak{N}_{s_{2}}(H)$ then one has
    $\operatorname{coeff}^{{}^{s_{0}}H}(\deg_{\mathcal{V}_{s_{1}m,j_{1}}}\cdot\deg_{\mathcal{V}_{s_{2}m,j_{2}}})=x_{0}^{2}|W(H)|({}^{s_{0}}H)$
    and the result follows from the fact that $x_{0}|W(H)|=2$. On the other hand, if either of $\dim\mathcal{V}_{m,j_{1}}^{H}$ and $\dim\mathcal{V}_{m,j_{2}}^{H}$ are even or if $\operatorname{gcd}(s_{1},s_{2})\neq s_{0}$, or if $s_{0}\notin\mathfrak{N}_{s_{1}}(H)\cap\mathfrak{N}_{s_{2}}(H)$ then the result follows immediately from the assumption that $s_{0}\notin\{s_{1},s_{2}\}$.
  2. $(ii)$
    Supposing instead that $s_{0}\in\{s_{1},s_{2}\}$, if it is also the case that $\dim\mathcal{V}_{m,j_{0}}^{H}$ is even, then one has $y_{m,j_{0}}=0$ and the result follows from (3). If one has $s_{0}\in\{s_{0},s_{1}\}$, with both $\dim\mathcal{V}_{m,j_{1}}^{H}$, $\dim\mathcal{V}_{m,j_{s_{2}}}^{H}$ odd and $s_{0}=\gcd(s_{1},s_{2})$ and $s_{0}\in\mathfrak{N}_{s_{1}}(H)\cap\mathfrak{N}_{s_{2}}(H)$, then (3) becomes
    $\deg_{\mathcal{V}_{s_{1}m,j_{1}}}\cdot\deg_{\mathcal{V}_{s_{2}m,j_{2}}}$ $=(G)-x_{0}({}^{s_{0}}H)+2x_{0}({}^{s_{0}}H)+a^{\prime},$
    where $a^{\prime}\in A(G)$ is such that $\operatorname{coeff}^{{}^{sm}H}(a^{\prime})=0$, and the result follows. On the other hand, if either of $\dim\mathcal{V}_{s_{1}m,j_{s_{1}}}^{H},\dim\mathcal{V}_{s_{1}m,j_{s_{1}}}^{H}$ are even (in which case one has $y_{m,j_{1}}y_{m,j_{2}}=0$) or if $\operatorname{gcd}(t,l)\neq s_{0}$, or if $s_{0}\notin\mathfrak{N}_{s_{1}}(H)\cap\mathfrak{N}_{s_{2}}(H)$, then the result follows from (3).
$\square$ What follows is the final generalization of the previous results to an orbit type of maximal kind $(H)\in\mathfrak{M}$ and its $s$-foldings $\{({}^{s}H)\}_{s\in{\mathbb{N}}}\subset\mathfrak{M}$ in the Burnside Ring product of any finite collection of basic degrees $\{\deg_{\mathcal{V}_{s_{k}m,j_{s_{k}}}}\}_{k=1}^{N}$ with $m>0$, $j_{s_{1}},\ldots,j_{s_{N}}\in\{0,1,\ldots,r\}$ and where $s_{1},\ldots,s_{N}\in\mathbb{N}$ are distinct.
Theorem 3.5.
Let $(H)\in\mathfrak{M}$ be an orbit type of maximal kind. For any finite collection of basic degrees $\{\deg_{\mathcal{V}_{s_{k}m,j_{s_{k}}}}\}_{k=1}^{N}$ with $m>0$, $j_{s_{1}},\ldots,j_{s_{N}}\in\{0,1,\ldots,r\}$ and where $s_{1},\ldots,s_{N}\in\mathbb{N}$ are distinct and for any number $s_{0}\in{\mathbb{N}}$, one has
$\operatorname{coeff}^{{}^{s_{0}}H}$ $\left(\prod\limits_{k=1}^{N}\deg_{\mathcal{V}_{s_{k}m,j_{s_{k}}}}\right)=-x_{0}\left[2\nmid\dim\mathcal{V}_{m,j_{s_{0}}}^{H}\right]\left[s_{0}\in\{s_{1},\ldots,s_{N}\}\right]$
$\quad+2x_{0}\sum\limits_{\begin{subarray}{c}I\subset\{s_{1},s_{2},\ldots,s_{N}\}\\ I\neq\emptyset,\{s_{0}\}\end{subarray}}(-2)^{|I|-2}\left[s_{0}\in\bigcap_{s\in I}\mathfrak{N}_{s}(H)\right]\left[s_{0}=\operatorname{gcd}(I)\right]\left[2\nmid\prod_{s\in I}\dim\mathcal{V}_{m,j_{s}}^{H}\right],$
where, if $s_{0}=s_{k}$ for any $k=1,\ldots,N$, then $j_{s_{0}}:=j_{s_{k}}$.
Proof.
Consider the Burnside Ring product of the relevant basic degrees
$\prod\limits_{k=1}^{N}\deg_{\mathcal{V}_{s_{k}m,j_{s_{k}}}}$ $=\prod\limits_{k=1}^{N}(G)-y_{m,j_{s_{k}}}({}^{s_{k}}H)+b_{m,j_{s_{k}}},$ (26)
and also the related Burnside Ring product
$\prod\limits_{k=1}^{N}(G)-y_{m,j_{s_{k}}}({}^{s_{k}}H).$ (27)
Notice that for any $s_{0}\in\mathbb{N}$, one has
$\operatorname{coeff}^{{}^{s_{0}}H}\left(\prod\limits_{k=1}^{N}\deg_{\mathcal{V}_{s_{k}m,j_{s_{k}}}}\right)=\operatorname{coeff}^{{}^{s_{0}}H}\left(\prod\limits_{k=1}^{N}(G)-y_{m,j_{s_{k}}}({}^{s_{k}}H)\right),$
and also that (27) has the expansion
$\prod\limits_{k=1}^{N}(G)-$ $y_{m,j_{s_{k}}}({}^{s_{k}}H)=\sum\limits_{I\subset\{s_{1},s_{2},\ldots,s_{N}\}}\prod\limits_{s\in I}-y_{m,j_{s}}({}^{s}H)$
$=\sum\limits_{I\subset\{s_{1},s_{2},\ldots,s_{N}\}}|W(H)|^{|I|-1}\left(\prod\limits_{s\in I}-y_{m,j_{s}}\right)\left[\gcd(I)\in\bigcap_{s\in I}\mathfrak{N}_{s}(H)\right]({}^{\operatorname{gcd}(I)}H),$
where the notation $I$ is used to indicate a subset of the indices $\{s_{1},s_{2},\ldots,s_{N}\}$ and the expression $\sum_{I\subset\{s_{1},s_{2},\ldots,s_{N}\}}$ describes a summation over all such subsets, including the empty set, in which case we put $\prod_{s\in\emptyset}-y_{m,j_{s}}(\alpha_{*})({}^{s}H):=(G)$, and the full set. It follows that the coefficient of $({}^{s_{0}}H)\in\Phi_{0}(G;\mathscr{H})$ in the Burnside Ring product (26), which for convenience of notation we denote by
$n_{s_{0}}:=\operatorname{coeff}^{{}^{s_{0}}H}(\prod\limits_{k=1}^{N}\deg_{\mathcal{V}_{s_{k}m,j_{s_{k}}}}),$
is specified by the formula
$n_{s_{0}}=\sum\limits_{\begin{subarray}{c}I\subset\{s_{1},s_{2},\ldots,s_{N}\}\\ I\neq\emptyset\end{subarray}}|W(H)|^{|I|-1}\left[s_{0}\in\bigcap_{s\in I}\mathfrak{N}_{s}(H)\right][s_{0}=\operatorname{gcd}(I)]\prod\limits_{s\in I}-y_{m,s}.$ (28)
Now, for any non-empty subset of indices $I\subset\{s_{1},s_{2},\ldots,s_{N}\}$, the product
$\prod\limits_{s\in I}-y_{m,j_{s}}=\prod\limits_{s\in I}-x_{0}\left[2\nmid\dim\mathcal{V}_{m,j_{s}}^{H}\right],$ (29)
is determined by the rule
$\prod\limits_{s\in I}-y_{m,j_{s}}=\begin{cases}(-x_{0})^{|I|}\quad&\text{ if }\dim\mathcal{V}_{m,j_{s}}^{H}\text{ is odd for all }s\in I;\\ 0\quad&\text{ otherwise}.\end{cases}$
Therefore, (28) becomes
$n_{s_{0}}$ $=\sum\limits_{\begin{subarray}{c}I\subset\{s_{1},s_{2},\ldots,s_{N}\}\\ I\neq\emptyset\end{subarray}}|W(H)|^{|I|-1}(-x_{0})^{|I|}\left[s_{0}\in\bigcap_{s\in I}\mathfrak{N}_{s}(H)\right]\left[s_{0}=\operatorname{gcd}(I)\right]\left[2\nmid\prod_{s\in I}\dim\mathcal{V}_{m,j_{s}}^{H}\right]$
$=-x_{0}\left[2\nmid\dim\mathcal{V}_{m,j_{0}}^{H}\right]\left[s_{0}\in\{s_{1},\ldots,s_{N}\}\right]$
$\quad+2x_{0}\sum\limits_{\begin{subarray}{c}I\subset\{s_{1},s_{2},\ldots,s_{N}\}\\ I\neq\emptyset,\{s_{0}\}\end{subarray}}(-2)^{|I|-2}\left[s_{0}\in\bigcap_{s\in I}\mathfrak{N}_{s}(H)\right]\left[s_{0}=\operatorname{gcd}(I)\right]\left[2\nmid\prod_{s\in I}\dim\mathcal{V}_{m,j_{s}}^{H}\right],$
where we have used the fact that $x_{0}|W(H)|=2$. $\square$ The following practical Corollary will prove useful in subsequent sections.
Corollary 3.6.
Let $(H)\in\mathfrak{M}$ be an orbit type of maximal kind. For any finite collection of basic degrees $\{\deg_{\mathcal{V}_{s_{k}m,j_{s_{k}}}}\}_{k=1}^{N}$ with $m>0$, $j_{s_{1}},\ldots,j_{s_{N}}\in\{0,1,\ldots,r\}$ and where $s_{1},\ldots,s_{N}\in\mathbb{N}$ are distinct and for any $s\in\{s_{1},\ldots,s_{N}\}$ with $2\nmid\dim\mathcal{V}_{m,j_{s}}^{H}$,
$\operatorname{coeff}^{{}^{s}H}$ $\left(\prod\limits_{k=1}^{N}\deg_{\mathcal{V}_{s_{k}m,j_{s_{k}}}}\right)\neq 0.$ (30)
Proof.
Indeed, notice that $-x_{0}+2x_{0}C\neq 0$ for any value of $C\in{\mathbb{N}}$. $\square$
4 On the Non-Triviality of the Degree $G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))$
Given an orbit type of maximal kind $(H)\in\mathfrak{M}$ and a positive number $s>0$, we put
$\begin{cases}\Sigma^{s}(H):=\{(m,j)\in\Sigma_{0}:2\nmid\dim\mathcal{V}_{m,j}^{{}^{s}H}\};\\ \mathfrak{n}^{s}(H):=|\Sigma^{s}(H)|,\end{cases}$
and
$S(H):=\{s\in{\mathbb{N}}:\mathfrak{n}^{s}(H)\neq 0\}.$
Remark 4.1.
If $(H)\in\mathfrak{M}_{m_{0}}$ for some $m_{0}>0$, then $(m,j)\in\Sigma^{s}(H)$ only if $m=m_{0}s$.
We are now in a position to present our main result.
Proposition 4.2.
Let $(H)\in\mathfrak{M}$ be an orbit type of maximal kind. For any $s_{0}\in{\mathbb{N}}$, one has
$\operatorname{coeff}^{{}^{s_{0}}H}$ $\left(G\text{\rm-deg}\left(\mathscr{A},B(\mathscr{H})\right)\right)=-x_{0}\left[2\nmid\mathfrak{n}^{s_{0}}(H)\right]\left[s_{0}\in S(H)\right]$
$\quad+2x_{0}\sum\limits_{\begin{subarray}{c}I\subset S(H)\\ I\neq\emptyset,\{s_{0}\}\end{subarray}}(-2)^{|I|-2}\left[s_{0}\in\bigcap_{s\in I}\mathfrak{N}_{s}(H)\right]\left[s_{0}=\operatorname{gcd}(I)\right]\left[2\nmid\prod_{s\in I}\mathfrak{n}^{s}(H)\right].$
Proof.
For convenience of notation, put
$\Sigma^{0}(H):=\{(m,j)\in\Sigma_{0}:2\mid\dim\mathcal{V}_{m,j}^{{}^{s}H}\text{for all }s>0\},$
and define
$\rho^{s}_{G}(H):=\prod_{(m,j)\in\Sigma^{s}(H)}\deg_{\mathcal{V}_{m,j}}.$
Enumerating the set $S(H)=\{s_{1},s_{2},\ldots,s_{N}\}$, one clearly has
$\Sigma_{0}\setminus\Sigma^{0}(H)=\bigcup_{k=1}^{N}\Sigma^{s_{k}}(H),$
such the degree computation (10) becomes
$G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))=\rho^{0}_{G}(H)\cdot\prod_{k=1}^{N}\rho^{s_{k}}_{G}(H).$
From Lemma 3.2 and Corollary 3.3, one has
$\rho_{G}^{s}(H)=\begin{cases}(G)+b_{0}&\text{ if }s=0;\\ (G)-y_{s}({}^{s}H)+b_{s}&\text{ if }s>0,\end{cases}$
where $b_{0},b_{s}$ satisfy $\operatorname{coeff}^{{}^{s}H}(b_{0})=\operatorname{coeff}^{{}^{s}H}(b_{s})=0$ for all $s\in{\mathbb{N}}$ and
$y_{s}=\begin{cases}0&\text{ if }\mathfrak{n}^{s}(H)\text{ is even};\\ x_{0}&\text{ if }\mathfrak{n}^{s}(H)\text{ is odd}.\end{cases}$
Assume, without loss of generality, that $(H)\in\mathfrak{M}_{m_{0}}$ and notice that each of the elements $\rho_{G}^{s_{k}}(H)$ has the form of a basic degree corresponding to an index $(s_{k}m_{0},j_{s_{k}})\in\Sigma$ satisfying
$2\mid\dim\mathcal{V}_{m_{0},j_{s_{k}}}^{H}+\mathfrak{n}^{s_{k}}(H),$
i.e., such that
$\operatorname{coeff}^{{}^{s_{k}}H}\left(\deg_{\mathcal{V}_{s_{k}m_{0},j_{s_{k}}}}\right)=\operatorname{coeff}^{{}^{s_{k}}H}\left(\rho_{G}^{s_{k}}(H)\right).$
Therefore, the coefficient of any $s_{0}$-folding $({}^{s_{0}}H)\in\mathfrak{M}$ in the degree (10) can be computed as follows
$\operatorname{coeff}^{{}^{s_{0}}H}(G\text{\rm-deg}(\mathscr{A},B(\mathscr{H})))=\operatorname{coeff}^{{}^{s_{0}}H}\left(\prod_{k=1}^{N}\deg_{\mathcal{V}_{s_{k}m_{0},j_{s_{k}}}}\right),$
and the result follows from Theorem 3.5. $\square$ Just as Corollary 3.6 naturally follows from Theorem 3.5, the following Corollary is a direct implication of Proposition 4.2.
Corollary 4.3.
Let $(H)\in\mathfrak{M}$ be an orbit type of maximal kind. For any $s\in{\mathbb{N}}$ with $2\nmid n^{s}(H)$, one has
$\operatorname{coeff}^{{}^{s}H}\left(G\text{\rm-deg}\left(\mathscr{A},B(\mathscr{H})\right)\right)\neq 0.$
In turn, combining Lemma 1.1 with Corollary 4.3 leads immediately to Theorem 1.6.
Appendix A The $G$-Equivariant Leray-Schauder Degree
Given an isometric Banach $G$-representation $\mathscr{H}$ of functions taking values in an orthogonal $\Gamma$-representation $V$, the $G$-equivariant Leray-Schauder Degree is a topological tool used to study the solution sets associated with equations of the form
$\mathscr{F}(u)=0,\quad u\in\mathscr{H},$ (31)
where $\mathscr{F}:\mathscr{H}\rightarrow\mathscr{H}$ is an operator satisfying the following conditions:
  1. ($B_{1}$)
    $\mathscr{F}$ is a $G$-equivariant completely continuous field;
  2. ($B_{2}$)
    there exists a sufficiently large $R>0$ such that $\mathscr{F}$ is $B_{R}(0)$-admissibly $G$-homotopic to the identity operator $\operatorname{Id}:\mathscr{H}\rightarrow\mathscr{H}$;
  3. ($B_{3}$)
    $D\mathscr{F}(0):\mathscr{H}\rightarrow\mathscr{H}$ exists and the operator $\operatorname{Id}-D\mathscr{F}(0):\mathscr{H}\rightarrow\mathscr{H}$ is a $G$-equivariant compact field;
  4. ($B_{4}$)
    if $D\mathscr{F}(0):\mathscr{H}\rightarrow\mathscr{H}$ is an isomorphism, there exists a sufficiently small $\epsilon>0$ such that $\mathscr{F}$ is $B_{\epsilon}(0)$-admissibly $G$-homotopic to $D\mathscr{F}(0)$.
Condition ($B_{1}$) allows the problem of the existence of solutions for equation (31) to be reformulated as a question concerning the non-triviality of the degree $G\text{\rm-deg}(\mathscr{F},B(\mathscr{H}))$. In turn, Conditions ($B_{2}$)-($B_{4}$) reduce calculation of $G\text{\rm-deg}(\mathscr{F},B(\mathscr{H}))$ to a Burnside Ring product involving a finite number of computationally simpler $G$-basic degrees. Equivariant notation. Let $G$ be a compact Lie group. For any subgroup $H\leq G$, we denote by $(H)$ its conjugacy class, by $N(H)$ its normalizer by $W(H):=N(H)/H$ its Weyl group in $G$. The set of all subgroup conjugacy classes in $G$ is denoted by $\Phi(G):=\{(H):H\leq G\}$ and has a natural partial order defined as follows
$(H)\leq(K)\iff\exists_{g\in G}\;\;gHg^{-1}\leq K.$
As is possible with any partially ordered set, we extend the natural order over $\Phi(G)$ to a total order, which we indicate by $<$ to differentiate the two relations. Moreover, we put $\Phi_{0}(G):=\{(H)\in\Phi(G):\text{$W(H)$ is finite}\}$ and, for any $(H),(K)\in\Phi_{0}(G)$, we denote by $n(H,K)$ the number of subgroups $\tilde{K}\leq G$ with $\tilde{K}\in(K)$ and $H\leq\tilde{K}$. Given a $G$-space $X$ with an element $x\in X$, we denote by $G_{x}:=\{g\in G:gx=x\}$ the isotropy group of $x$ and we call $(G_{x})\in\Phi(G)$ the orbit type of $x\in X$. Put $\Phi(G,X):=\{(H)\in\Phi_{0}(G):(H)=(G_{x})\;\text{for some $x\in X$}\}$ and $\Phi_{0}(G,X):=\Phi(G,X)\cap\Phi_{0}(G)$. For a subgroup $H\leq G$, the subspace $X^{H}:=\{x\in X:G_{x}\geq H\}$ is called the $H$-fixed-point subspace of $X$. If $Y$ is another $G$-space, then a continuous map $f:X\to Y$ is said to be $G$-equivariant if $f(gx)=gf(x)$ for each $x\in X$ and $g\in G$. The Burnside Ring. The free $\mathbb{Z}$-module $A(G):=\mathbb{Z}[\Phi_{0}(G)]$ has a natural ring structure when equipped with the multiplicative operation defined, for any pair of generators $(H),(K)\in\Phi_{0}(G)$, as follows
$(H)\cdot(K):=\sum\limits_{(L)\in\Phi_{0}(G)}n_{L}(L),$
and where the coefficients $n_{L}\in\mathbb{Z}$ are given by the recurrence formula
$n_{L}:=\frac{n(L,H)|W(H)|n(L,K)|W(K)|-\sum_{(\tilde{L})>(L)}n_{\tilde{L}}n(L,\tilde{L})|W(\tilde{L})|}{|W(L)|}.$
Any Burnside Ring element $a\in A(G)$ can be expressed as a formal sum over some finite number of generator elements
$a=n_{1}(H_{1})+n_{2}(H_{2})+\cdots+n_{N}(H_{N}),$
and we use the notation
$\operatorname{coeff}^{H}(a)=n_{H},$
to specify the integer coefficient standing next to the generator element $(H)\in\Phi_{0}(G)$. An Axiomatic Construction of the Leray-Schauder $G$-Equivariant Degree. Let $\mathscr{E}$ be any isometric Banach $G$-representation. A map $f:\mathscr{E}\rightarrow\mathscr{E}$ is said to be a completely continuous $G$-equivariant field if it can be expressed in the form
$f(x)=x-F(x),$
for some compact $G$-equivariant map $F:\mathscr{E}\rightarrow\mathscr{E}$. Moreover, given an open bounded $G$-invariant set $\Omega\subset\mathscr{E}$, a completely continuous $G$-equivariant field is said to be $\Omega$-admissible and the pair $(f,\Omega)$ is called an admissible $G$-pair if $f(x)\neq 0$ for all $x\in\partial\Omega$. We denote by $\mathscr{M}^{G}(\mathscr{E})$ the set of all admissible $G$-pairs in $\mathscr{E}$ and by $\mathscr{M}^{G}$ the set of all admissible $G$-pairs defined by taking a union over all isometric Banach $G$-representations as follows
$\mathscr{M}^{G}:=\bigcup_{\mathscr{E}}\mathscr{M}^{G}(\mathscr{E}).$
The $G$-equivariant Leray-Schauder degree is defined as the unique map $G\text{\rm-deg}:\mathscr{M}^{G}\rightarrow A(G)$ that assigns to every admissible $G$-pair $(f,\Omega)$ the Burnside Ring element
$G\text{\rm-deg}(f,\Omega)=\sum_{(H)\in\Phi_{0}(G)}{n_{H}(H)},$ (32)
satisfying the four degree axioms
  • (Existence) If $n_{H}\neq 0$ for some $(H)\in\Phi_{0}(G)$ in (32), then there exists $x\in\Omega$ such that $f(x)=0$ and $(G_{x})\geq(H)$.
  • (Additivity) For any two disjoint open $G$-invariant subsets $\Omega_{1}$ and $\Omega_{2}$ with $f^{-1}(0)\cap\Omega\subset\Omega_{1}\cup\Omega_{2}$, one has
    $G\text{\rm-deg}(f,\Omega)=G\text{\rm-deg}(f,\Omega_{1})+G\text{\rm-deg}(f,\Omega_{2}).$
  • (Homotopy) For any $\Omega$-admissible $G$-homotopy, $h:[0,1]\times V\to V$, one has
    $G\text{\rm-deg}(h_{t},\Omega)=\mathrm{constant}.$
  • (Normalization) For any open bounded neighborhood of the origin in an isometric Banach $G$-representation $\mathscr{E}$ with the identity operator $\operatorname{Id}:\mathscr{E}\rightarrow\mathscr{E}$, one has
    $G\text{\rm-deg}(\operatorname{Id},\Omega)=(G).$
[…]
  • (Multiplicativity) For any $(f_{1},\Omega_{1}),(f_{2},\Omega_{2})\in\mathcal{M}^{G}$,
    $G\text{\rm-deg}(f_{1}\times f_{2},\Omega_{1}\times\Omega_{2})=G\text{\rm-deg}(f_{1},\Omega_{1})\cdot G\text{\rm-deg}(f_{2},\Omega_{2}),$
    where the multiplication $\cdot$’ is taken in the Burnside ring $A(G)$.
  • (Recurrence Formula) For an admissible $G$-pair $(f,\Omega)\in\mathscr{M}^{G}(\mathscr{E})$, the $G$-degree (32) can be computed using the following Recurrence Formula:
    $n_{H}=\frac{\deg(f^{H},\Omega^{H})-\sum_{(K)>(H)}{n_{K}\,n(H,K)\,\left|W(K)\right|}}{\left|W(H)\right|},$ (33)
    where $\left|X\right|$ stands for the number of elements in the set $X$ and $\deg(f^{H},\Omega^{H})$ is the Brouwer degree of the map $f^{H}:=f|_{\mathscr{E}^{H}}$ on the set $\Omega^{H}\subset\mathscr{E}^{H}$.
Computation of Brouwer equivariant degree. Let’s denote by $\{\mathcal{U}_{i}\}_{i\in\mathbb{N}}$ the set of all irreducible $G$-representations and define the $i$-th basic degree as follows
$\deg_{\mathcal{U}_{i}}:=G\text{\rm-deg}(-\operatorname{Id},B(\mathcal{U}_{i})).$
Given any isometric Banach $G$-representation with a $G$-isotypic decomposition
$\mathscr{E}=\bigoplus_{i\in\mathbb{N}}\mathscr{E}_{i},$
and any bounded $G$-equivariant linear isomorphism $T:\mathscr{E}\to\mathscr{E}$, the Multiplicativity and Homotopy properties of the $G$-equivariant Brouwer degree, together with Schur’s Lemma implies
$G\text{\rm-deg}(\operatorname{Id}-T,B(\mathscr{E}))=\prod_{i\in\mathbb{N}}G\text{\rm-deg}(T_{i},B(\mathscr{E}_{i}))=\prod_{i\in\mathbb{N}}\prod_{\mu\in\sigma_{-}(T)}\left(\deg_{\mathcal{U}_{i}}\right)^{m_{i}(\mu)}$
where $T_{i}=\operatorname{Id}-T|_{\mathscr{E}_{i}}:\mathscr{E}_{i}\rightarrow\mathscr{E}_{i}$ and $\sigma_{-}(T)$ denotes the real negative spectrum of $T$.
Notice that each of the basic degrees:
$\deg_{\mathcal{U}_{i}}=\sum_{(H)\in\Phi_{0}(G)}n_{H}(H),$
can be practically computed, using the recurrence formula (33), as follows
$n_{H}=\frac{(-1)^{\dim\mathcal{U}_{i}^{H}}-\sum_{H<K}{n_{K}\,n(H,K)\,\left|W(K)\right|}}{\left|W(H)\right|}.$
The following fact is well-known (see for example [1]).
Lemma A.1.
For any irreducible $G$-representation $\mathcal{U}$, the basic degree $\deg_{\mathcal{U}}\in A(G)$ is an involutive element of the Burnside Ring, i.e.
$(\deg_{\mathcal{U}})^{2}=\deg_{\mathcal{U}}\cdot\deg_{\mathcal{U}}=(G).$