### v3

# Patterns of Non-Stationary Solutions to Symmetric Systems of Second-Order Differential Equations

Abstract
We establish the existence of non-stationary solutions to a symmetric system of second-order autonomous differential equations. Our technique is based on the equivariant degree theory and involves a novel characterization of orbit types of maximal kind in the Burnside Ring product of a finite number of basic degrees for the group $O(2)\times\Gamma\times\mathbb{Z}_{2}$.
Mathematics Subject Classification: Primary: 34C25, 37C81, 47H11, 55M25, 19A22
Key Words and Phrases: symmetric equation, existence of solutions, equivariant Leray-Schauder degree, nonlinear analysis, periodic solutions.
1 Introduction
The classical forced pendulum equation $\ddot{x}=\bm{f}(x)$ serves as a paradigmatic problem for validating novel techniques in nonlinear analysis (cf. [20]-[24]). More generally, dynamical systems described by second-order differential equations strike a delicate balance between simplicity and a capacity for capturing rich and intricate nonlinear behavior, making them ideal candidates for the testing of analytical methods. On the other hand, autonomous differential equations with periodic boundary conditions always admit natural $SO(2)$ symmetries related to time shifting that make them particularly well-suited for a framework based on the equivariant degree theory. Second order autonomous differential equations not involving any first derivative terms, specifically, admit $O(2)$ symmetries, where the reflection generator acts by time reversal. When these equations are subjected to an odd forcing function and arranged in symmetric networks characterized by a finite group $\Gamma$, their full symmetry group becomes
$G:=O(2)\times\Gamma\times{\mathbb{Z}}_{2}.$
Our primary objective is to demonstrate how the $G$-equivariant Leray-Schauder degree can be used to establish the existence of non-stationary solutions to the following $\Gamma$-symmetric system of $p$-periodic, second order, autonomous equations:
$\begin{cases}\ddot{x}(t)=f(x(t))+Ax(t),\quad t\in\mathbb{R},\;x(t)\in V\\ x(t)=x(t+p),\;\dot{x}(t)=\dot{x}(t+p),\end{cases}$ (1)
where $V:=\mathbb{R}^{N}$ is an orthogonal $\Gamma$-representation, $A:V\rightarrow V$ is a symmetric $\Gamma$-equivariant matrix and $f:V\rightarrow V$ is a continuous function satisfying the conditions
  1. ($A_{1}$)
    $f$ is $\Gamma$-equivariant, i.e. $f(\gamma x)=\gamma f(x)$ for all $\gamma\in\Gamma,\;t\in\mathbb{R},\;x\in V$.
  2. ($A_{2}$)
    $f$ is odd, i.e. $f(-x)=-f(x)$ for all $x\in V$.
  3. ($A_{3}$)
    $f(x)$ is $o(|x|)$ as $x$ approaches 00, i.e.
    $\lim_{x\rightarrow 0}\frac{f(x)}{|x|}=0.$
  4. ($A_{4}$)
    there exists a constant $M>0$ such that for all $x\in V$ with $|x|\geq M$, one has
    $f(x)\cdot x>0.$
Conditions ($A_{1}$)($A_{2}$) ensure that the equations admit the symmetries of the product group $O(2)\times\Gamma\times{\mathbb{Z}}_{2}$, while conditions ($A_{3}$) and ($A_{4}$) imply the well-behavedness of our non-linearity at solutions to (1) near zero and infinity, respectively. Specifically, ($A_{2}$) allows the equivariant degree to distinguish between stationary and non-stationary functions, ($A_{3}$) guarantees that the nonlinearity does not interact with the spectrum of $A$ and the Nagumo condition ($A_{4}$) is essential for establishing a-priori bounds on the magnitude of solutions to (1). System (1) is a natural model for the dynamics of a configuration of $N$ identical pendula, subjected to nonlinear forcing $f$ and with coupling relations specified by the matrix $A$ and the group $\Gamma$. In this context, the vector $x(t):=(x_{1}(t),x_{2}(t),\ldots,x_{N}(t))$ represents the angles of the pendula from the vertical position, $\dot{x}(t):=(\dot{x}_{1}(t),\dot{x}_{2}(t),\ldots,\dot{x}_{N}(t))$ the angular velocities and $\ddot{x}(t):=(\ddot{x}_{1}(t),\ddot{x}_{2}(t),\ldots,\ddot{x}_{N}(t))$ the angular accelerations at any time $t>0$. Methods based on the Leray-Schauder degree have been used to solve a wide variety of second order differential equations [3, 9, 14, 17, 18, 22, 24]. However, the inability of the Leray-Schauder degree to distinguish between periodic solutions that differ by a fixed time shift and differentiate constant solutions from non-constant solutions have hindered its suitability for autonomous differential equations. The effectiveness of the $G$-equivariant Leray-Schauder degree in detecting non-constant solutions to the system (1) for several explicit choices of the symmetry group $\Gamma$ has already been demonstrated by Garcia-Azpetia et al. in the article [13]. More generally, the $G$-equivariant Leray-Schauder degree has been applied to problems admitting appropriate spatial-temporal symmetries [4, 5, 11, 12, 15, 19]. Compared with other techniques to solve systems similar to (1), the equivariant degree approach requires relatively minimal regularity assumptions on the nonlinearity. For example, existence results were obtained in [2] by topological index argument and in [10] using Morse theory framework, under the assumption that the right-hand-side interacts with only finitely many eigenvalues of the differential operator. A prerequisite for the application of any $G$-equivariant degree-theory based argument to (1) is the reformulation of our system as a fixed point equation in an appropriate functional $G$-space with a nonlinear operator in the form of a $G$-equivariant compact perturbation of identity. In [13], Garcia-Azpetia et al. prove that the problem of finding $p$-periodic solutions to (1) is equivalent to the problem of finding zeros of the operator
$\mathscr{F}:\mathscr{H}\rightarrow\mathscr{H},\quad\mathscr{F}(u):=u-\mathscr{L}^{-1}(\beta^{2}f(u)+\beta^{2}Au-u),$
where $\beta:=\frac{p}{2\pi}$, $\mathscr{H}:=H^{2}_{2\pi}([0,2\pi];V)$ and $\mathscr{L}:\mathscr{H}\rightarrow\mathscr{H}$ is the differential operator $\mathscr{L}u:=\ddot{u}-u$. Under the conditions ($A_{1}$)($A_{4}$) and provided that the linearization $D\mathscr{F}(0):\mathscr{H}\rightarrow\mathscr{H}$ is an isomorphism, Garcia-Azpetia et al. additionally demonstrate that:
  1. $\rm(i)$
    $\mathscr{F}$ is a $G$-equivariant completely continuous field with respect to the natural isometric action of $G$ on $\mathscr{H}$;
  2. $\rm(ii)$
    all of the non-trivial solutions to (1) can be confined to an annulus of the form $\Omega:=B_{R}(\mathscr{H})\setminus B_{\varepsilon}(\mathscr{H})\subset\mathscr{H}$ for some sufficiently small choice of $\varepsilon>0$ and sufficiently large choice of $R>0$;
  3. $\rm(iii)$
    $(\mathscr{F},\Omega)$ constitutes an admissible $G$-pair, such that the existence of non-trivial solutions to (1) is equivalent to non-triviality of the degree invariant
    $G\text{\rm-deg}(\mathscr{F},\Omega)=(G)-G\text{\rm-deg}(D\mathscr{F}(0),B_{1}(\mathscr{H})),$
[…]
Together with a direct application of the existence property of the $G$-equivariant Leray-Schauder degree (cf. Appendix A), the above considerations suggest the following sufficient condition for the existence of a non-trivial solution to (1):
Lemma 1.1.
If for some orbit type $(H)\in\Phi_{0}(G)\setminus\{(G)\}$ one has
$\operatorname{coeff}^{H}\left(G\text{\rm-deg}(D\mathscr{F}(0),B(\mathscr{H}))\right)\neq 0,$
(cf. Appendix A for definition of the coefficient operator $\operatorname{coeff}^{H}:A(G)\rightarrow{\mathbb{Z}}$) then there exists a function $u\in\mathscr{H}\setminus\{0\}$ satisfying (4) with an isotropy subgroup $G_{u}\leq G$ satisfying $(G_{u})\geq(H)$.
In this way, the problem of finding $p$-periodic solutions to (1) has been reformulated as the problem of computing the coefficients of non-unit orbit types (i.e. orbit types $(H)\in\Phi_{0}(G)\setminus\{(G)\}$) in the Burnside ring element $G\text{\rm-deg}(D\mathscr{F}(0),B(\mathscr{H}))$. Moreover, since the $G$-equivariant Leray-Schauder degree provides a full equivariant classification to the solution set of $\mathscr{F}(u)=0,\;u\in\mathscr{H}$, Lemma 1.1 represents a framework for identifying the exact spatio-temporal symmetries of any predicted solution. In order to take advantage of the full power of the $G$-equivariant Leray-Schauder degree for this purpose, we must first identify the irreducible representations of $G$ and describe the $G$-isotypic decomposition of $\mathscr{H}$. Without specifying the group $\Gamma$, we assume that a complete list of the irreducible $\Gamma$-representations $\{\mathcal{V}_{j}\}_{j=0}^{r}$ is made available. As a $\Gamma$-representation, the space $V$ is also a natural $\Gamma\times\mathbb{Z}_{2}$-representation with the $\Gamma\times\mathbb{Z}_{2}$-isotypic decomposition
$V=V_{0}\oplus V_{1}\oplus\cdots\oplus V_{r},$
where each $\Gamma\times\mathbb{Z}_{2}$-isotypic component $V_{j}$ is modeled on the irreducible $\Gamma\times\mathbb{Z}_{2}$-representation $\mathcal{V}_{j}^{-}$ (here the superscript is meant to indicate that $\mathcal{V}_{j}$ has been equipped with antipodal $\mathbb{Z}_{2}$-action) in the sense that $V_{j}$ is equivalent to the direct sum of some finite number of copies of $\mathcal{V}_{j}^{-}$, i.e.
$V_{j}\simeq\mathcal{V}_{j}^{-}\oplus\cdots\oplus\mathcal{V}_{j}^{-}.$
The exact number of irreducible $\Gamma$-representations $\mathcal{V}_{j}^{-}$ ‘contained’ in the $\Gamma\times\mathbb{Z}_{2}$-isotypic component $V_{j}$ is called the $\mathcal{V}_{j}$-isotypic multiplicity of $V$ and is calculated according to the ratio:
$m_{j}:=\dim V_{j}/\dim\mathcal{V}_{j}^{-},\quad j\in\{0,1,\ldots,r\}.$ (2)
Since $A:V\rightarrow V$ is $\Gamma$-equivariant, one has
$\sigma(A)=\bigcup_{j=0}^{r}\sigma(A_{j}),\quad A_{j}:=A|_{V_{j}}:V_{j}\rightarrow V_{j}.$
For the sake of simplicity, we choose to sidestep a fixed point argument made by Garcia-Azpetia et al. in [13] by enforcing the following non-degeneracy assumption:
  1. ($A_{0}$)
    for each $\Gamma$-isotypic component $V_{j}$, there exists a number $\mu_{j}\in\mathbb{R}$ with $\mu_{j}\neq-m^{2}/\beta^{2}$ for all $m\in{\mathbb{N}}$ such that
    $A_{j}=\mu_{j}\operatorname{Id}|_{V_{j}}:V_{j}\rightarrow V_{j}.$
On the other hand, for each $m\in{\mathbb{N}}$, we denote by $\mathcal{W}_{m}\simeq\mathbb{C}$ the irreducible $O(2)$-representation equipped with the $m$-folding $O(2)$-action
$e^{i\theta}w:=e^{im\theta}w,\quad\kappa w:=\bar{w},\quad e^{i\theta},\kappa\in O(2),\;w\in\mathcal{W}_{m},$
(here, $\kappa$ indicates the reflection generator in $O(2)$) and by $\mathcal{W}_{0}\simeq\mathbb{R}$ the irreducible $O(2)$-representation on which $O(2)$ acts trivially such that the space $\mathscr{H}$ admits the $O(2)$-isotypic decomposition
$\mathscr{H}:=\overline{\bigoplus\limits_{m=0}^{\infty}\mathscr{H}_{m}},\quad\mathscr{H}_{m}:=\{u\in\mathscr{H}:u(t)=\cos(mt)a+\sin(mt)b,\;a,b\in V\},$
where
$\mathscr{H}_{m}\simeq\mathcal{W}_{m}\otimes V,\quad m\in{\mathbb{N}}\cup\{0\}.$
The $G$-isotypic decomposition of $\mathscr{H}$ can now be expressed in terms of $G$-isotypic components $\mathscr{H}_{m,j}\subset\mathscr{H}$ modeled on the irreducible $G$-representations
$\mathcal{V}_{m,j}:=\mathcal{W}_{m}\otimes\mathcal{V}_{j}^{-},\quad m\in{\mathbb{N}}\cup\{0\},\;j\in\{0,1,\ldots,r\},$
as follows
$\mathscr{H}=\bigoplus_{j=0}^{r}\overline{\bigoplus\limits_{m=0}^{\infty}\mathscr{H}_{m,j}},$ (3)
where
$\mathscr{H}_{m,j}:=\{u\in\mathscr{H}:u(t)=\cos(mt)a+\sin(mt)b,\;a,b\in V_{j}\}.$
We are now in a position to present our main existence results:
Theorem 1.2.
Take $m>0$ and let $(H)$ be a maximal element in the isotropy lattice $\Phi_{0}(G;\mathscr{H}_{m}\setminus\{0\})$. Under the assumptions ($A_{0}$)($A_{4}$), if the matrix $A:V\rightarrow V$ admits an odd number of eigenvalues $\mu_{j}\in\sigma(A)$ with odd isotypic multiplicity $m_{j}$ (cf. (2)) satisfying
$\mu_{j}<-\left(\frac{m}{\beta}\right)^{2}\text{ and }2\nmid\dim\mathcal{V}_{m,j}^{H},$
then there exists a non-stationary solution $u\in\mathscr{H}\setminus\{0\}$ to the system (1) with an isotropy subgroup $G_{u}\leq G$ satisfying $(G_{u})\geq(H)$.
Remark 1.3.
In the field of equivariant bifurcation theory, the Ize Conjecture (IC) refers to the proposition that every irreducible representation of a compact Lie group admits at least one subgroup with an odd dimensional fixed point space. Although the IC was shown to be false by Lauterbach et al. in [25] via counterexample, the conjecture holds in many cases. For example, in Section 7.1, we use G.A.P. to numerically verify that the irreducible representations of the group $G=O(2)\times\Gamma\times{\mathbb{Z}}_{2}$, with the assignment $\Gamma=D_{8}$, satisfy the IC.
The remainder of this paper is organized as follows: In Section 2, we prepare the problem (1) for application of the $G$-equivariant Leray-Schauder degree. This involves a functional reformulation in an appropriate Sobolev space, allowing us to derive a practical formula for a degree invariant, in terms of the Burnside ring product of a finite number of basic degrees defined on the irreducible $G$-representations, whose non-triviality is equivalent to the existence of non-trivial solutions to (1). Using this framework, we analyze the maximal elements in the isotropy lattices $\Phi_{0}(G;\mathscr{H}_{m,j}\setminus\{0\})$, called orbit types of maximal kind in $\Phi_{0}(G)$. The significance of these orbit types is two-fold: first, the coefficients corresponding to orbit types of maximal kind in the degree invariant are computationally accessible and second, the non-triviality of these coefficients guarantees the existence of non-stationary solutions to (1). In Section 4, we present a classification for the generators of the Burnside ring $A(G)$ according to their lattice relations with respect to the $s$-folding homomorphism (13) and employ this classification, in Section 5, to derive a practical formula for characterizing the non-triviality of an orbit type of maximal kind in the product of a finite number of basic degrees. Specifically, we find that the contribution of each eigenvalue $\mu_{m,j}\in\sigma(D\mathscr{F}(0))$ to the appearance of an orbit type of maximal kind $(H)$ in the degree invariant $G\text{\rm-deg}(\mathscr{F},\Omega)$ depends on $(\rm i)$ the sign of $\mu_{m,j}$ and $(\rm ii)$ the parity of the dimension of the corresponding $H$-fixed-point space $\mathcal{V}_{m,j}^{H}$. Finally, in Section 7, we consider the implications of our main result for dihedral symmetry group $\Gamma=D_{N}$. For the special case of $D_{8}$ coupling symmetries, we use GAP to numerically verify $(\rm i)$ the pairwise disjointness of the maximal orbit type sets in each isotropy lattice $\Phi_{0}(G;\mathscr{H}_{m,j}\setminus\{0\})$ and $(\rm ii)$ the odd-dimensionality of the fixed point sets $\mathcal{V}_{m,j}^{H}$ for every maximal element in $\Phi_{0}(G;\mathscr{H}_{m,j}\setminus\{0\})$, allowing us to reformulate Theorem 1.2 in Section 7.1 as follows:
Proposition 1.4.
Let $\Gamma=D_{8}$ and take any fixed $m>0$. Under the assumptions ($A_{0}$)($A_{4}$), if the $j$-th eigenvalue of the matrix $A$ satisfies
$\mu_{j}<-\left(\frac{m}{\beta}\right)^{2},$
then, corresponding to every maximal element $(H)$ in the isotropy lattice $\Phi_{0}(G;\mathscr{H}_{m,j}\setminus\{0\})$, there exists a non-stationary solution $u\in\mathscr{H}\setminus\{0\}$ to the system (1) with an isotropy subgroup $G_{u}\leq G$ satisfying $(G_{u})\geq(H)$.
For convenience, the Appendices include an explanation of notations used, and a brief introduction to the $G$-equivariant Leray-Schauder degree. Readers who are interested in a deeper exposition of these topics are referred to [7, 8].
2 Setting for the Leray-Schauder $G$-Equivariant Degree
Notice that, with the substitutions $\beta:=p/2\pi$ and $u(t):=x(\beta t)$, the system (1) becomes
$\begin{cases}\ddot{u}(t)=\beta^{2}f(u(t))+\beta^{2}Au(t),\quad t\in\mathbb{R},\;u(t)\in V\\ u(t)=u(t+2\pi),\;\dot{u}(t)=\dot{u}(t+2\pi).\end{cases}$ (4)
Clearly, any $2\pi$-periodic solution $u(t):\mathbb{R}\rightarrow V$ for the system (4) corresponds to the $p$-periodic solution $x(t)=u(t/\beta)$ for our original equation. Following [13], we denote by $\mathscr{H}$ the Sobolev space of $2\pi$-periodic, $V$-valued functions
:=$\mathscr{H}:=\{u:\mathbb{R}\rightarrow V:u(0)=u(2\pi),\;u|_{[0,2\pi]}\in H^{2}(S^{1};V)\},$
equipped with the standard inner product
[…]
and the isometric $G$-action
$(e^{i\theta},\gamma,\pm 1)u(t):=\pm\gamma u(t+\theta),\quad(\kappa,\gamma,\pm 1)u(t):=\pm\gamma u(-t),\quad u\in\mathscr{H},\;t\geq 0,$
where $(e^{i\theta},\gamma,\pm 1),(\kappa,\gamma,\pm 1)\in O(2)\times\Gamma\times{\mathbb{Z}}_{2}$. We also consider the differential operator
$\mathscr{L}:\mathscr{H}\rightarrow\mathscr{E},\quad\mathscr{L}u:=\ddot{u}-u,$
the Nemytskii operator
$N:L^{2}(S^{1};V)\rightarrow L^{2}(S^{1};V),\quad(Nu)(t):=f(u(t)),$
and the Banach embedding
$j:\mathscr{H}\rightarrow L^{2}(S^{1};V),\quad j(u)(t):=u(t).$
Since $\mathscr{L}:\mathscr{H}\rightarrow L^{2}(S^{1};V)$ is a linear isomorphism, $N:L^{2}(S^{1};V)\rightarrow L^{2}(S^{1};V)$ is continuous and $j:\mathscr{H}\rightarrow L^{2}(S^{1};V)$ is compact, the operator $\mathscr{F}:\mathscr{H}\rightarrow\mathscr{H}$ given by
$\mathscr{F}(u):=u-\mathscr{L}^{-1}(\beta^{2}N(j(u))+\beta^{2}Aj(u)-j(u)),$ (5)
is a compact perturbation of the identity on $\mathscr{H}$ and the system (4) is equivalent to the operator equation
$\mathscr{F}(u)=0,$ (6)
in the sense that a function $u\in\mathscr{H}$ is a solution to (4) if and only if it satisfies (6). In what follows, (4) will be called the operator equation associated with the system (4). Gracia-Azpetia et al. demonstrate the applicability of the $G$-equivariant Leray-Schauder degree to (6) by proving that the assumptions ($A_{1}$)($A_{4}$) imply that the nonlinear operator (5) is a completely continuous $G$-equivariant field, differentiable at the origin $0\in\mathscr{H}$ with
$\mathscr{A}u:=\operatorname{Id}-\mathscr{L}^{-1}(\beta^{2}Au-u),\quad\mathscr{A}:=D\mathscr{F}(0):\mathscr{H}\rightarrow\mathscr{H}.$
2.1 Towards a Computational Formula for $G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))$
In order to effectively make use Lemma 1.1 to determine the existence of non-trivial solutions to (1), we must derive a practical formula for the computation of the local bifurcation invariant $G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))\in A(G)$. Our first step in this direction will be to take advantage of the $G$-isotypic decomposition (3) to collect spectral data related to the $G$-equivariant linear operator $\mathscr{A}:\mathscr{H}\rightarrow\mathscr{H}$. By Schur’s Lemma, one has
$\mathscr{A}(\mathscr{H}_{m,j})\subset\mathscr{H}_{m,j},\quad m\in{\mathbb{N}}\cup\{0\},\;j\in\{0,1,\ldots,r\},$
such that, adopting the notation
$\mathscr{A}_{m,j}:=\mathscr{A}|_{\mathscr{H}_{m,j}}:\mathscr{H}_{m,j}\rightarrow\mathscr{H}_{m,j},$
the spectrum of $\mathscr{A}$ admits the following decomposition
$\sigma(\mathscr{A})=\bigcup\limits_{m=0}^{\infty}\bigcup\limits_{j=0}^{r}\sigma(\mathscr{A}_{m,j}).$
Under assumption ($A_{0}$), the spectrum of each matrix $\mathscr{A}_{m,j}$ consists of the nonzero eigenvalue
$\mu_{m,j}:=1+\frac{\beta^{2}\mu_{j}-1}{1+m^{2}}=\frac{m^{2}+\beta^{2}\mu_{j}}{1+m^{2}},$ (7)
with multiplicity $m_{j}\in{\mathbb{N}}$ (cf. discussion of isotypic multiplicity (2)). The product property of the $G$-equivariant Leray-Schauder degree (cf. Appendix A) permits us to express the degree $G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))$ in terms of a Burnside ring product of the $G$-equivariant Leray-Schauder degrees of the various restrictions $\mathscr{A}_{m,j}:\mathscr{H}_{m,j}\rightarrow\mathscr{H}_{m,j}$ of the $G$-equivariant linear isomorphism $\mathscr{A}:\mathscr{H}\rightarrow\mathscr{H}$ to the $G$-isotypic components $\mathscr{H}_{m,j}$ on their respective open unit balls $B(\mathscr{H}_{m,j}):=\{u\in\mathscr{H}_{m,j}\;:\;\|u\|_{\mathscr{H}}<1\}$ as follows
$G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))=\prod\limits_{j=0}^{r}\prod\limits_{m=0}^{\infty}G\text{\rm-deg}(\mathscr{A}_{m,j},B(\mathscr{H}_{m,j})).$ (8)
On the other hand, each degree $G\text{\rm-deg}(\mathscr{A}_{m,j},B(\mathscr{H}_{m,j}))$ is fully specified by the $\mathcal{V}_{j}$-isotypic multiplicities $\{m_{j}\}_{j=0}^{r}$ together with the real spectra of $\mathscr{A}_{m,j}$ according to formula:
$G\text{\rm-deg}(\mathscr{A}_{m,j},B(\mathscr{H}_{m,j}))=\begin{cases}(\deg_{\mathcal{V}_{m,j}})^{m_{j}}\quad&\text{ if }\mu_{m,j}<0;\\ (G)\quad&\text{ otherwise,}\end{cases}$ (9)
where $\deg_{\mathcal{V}_{m,j}}\in A(G)$ is the basic degree associated with the irreducible $G$-representation $\mathcal{V}_{m,j}$ (cf. Appendix A) and $(G)\in A(G)$ is the unit element of the Burnside Ring. In addition, since each basic degree $\deg_{\mathcal{V}_{m,j}}$ is involutive in the Burnside ring (cf. Appendix A), one has
$(\deg_{\mathcal{V}_{m,j}})^{m_{j}}=\begin{cases}\deg_{\mathcal{V}_{m,j}}\quad&\text{ if }2\nmid m_{j};\\ (G)\quad&\text{ otherwise.}\end{cases}$ (10)
Putting together (8), (9) and (10), we introduce some notations to keep track of the indices
$\Sigma:=\left\{(m,j):m\in\mathbb{N}\cup\{0\},\;j\in\{0,1,\ldots,r\}\right\},$
which contribute non-trivially to $G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))$. To begin, the negative spectrum of $\mathscr{A}:\mathscr{H}\rightarrow\mathscr{H}$ is accounted for with the index set
$\Sigma_{-}:=\left\{(m,j)\in\Sigma:\mu_{m,j}<0\right\}.$
Combining this with formulas (8) and (9) yields
$G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))\;=\prod\limits_{(m,j)\in\Sigma_{-}}(\deg_{\mathcal{V}_{m,j}})^{m_{j}}.$
This computation can be further reduced by discarding the indices $(m,j)\in\Sigma$ associated with even $\mathcal{V}_{j}$-isotypic multiplicities $m_{j}$, whose corresponding basic degrees contribute trivially to the Burnside product (8). Specifically, we put
[…]
which, together with (10), permits us to express $G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))$ as the Burnside ring product of a finite number of unique basic degrees, as follows
$G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))\;=\prod\limits_{(m,j)\in\Sigma_{0}}\deg_{\mathcal{V}_{m,j}}.$ (11)
3 Orbit Types of Maximal Kind
It is always possible to study the non-triviality of the product of a finite number of Burnside ring elements (such as the product of basic degrees in the expression (11)) as a problem concerning the non-triviality of the product of a finite number of generator pairs $(K),(L)\in\Phi_{0}(G)$. Keeping in mind that the non-triviality of
$(K)\cdot(L)\in A(G),$
is equivalent to the existence of an orbit type $(H)\in\Phi_{0}(G)\setminus\{(G)\}$ with $\operatorname{coeff}^{H}((K)\cdot(L))\neq 0$, notice that the recurrence formula for the Burnside ring product (cf. Appendix A)
$C^{H}(K,L)=$
$\quad\frac{n(H,L)|W(L)|n(H,K)|W(K)|-\sum_{(\tilde{H})>(H)}C^{\tilde{H}}(K,L)n(H,\tilde{H})|W(\tilde{H})|}{|W(H)|},$
(here, we have temporarily adopted the notation $C^{H}(K,L):=\operatorname{coeff}^{H}((K)\cdot(L))$ due to constraints on space) is only useful if one first identifies the set of all orbit types $(\tilde{H})\in\Phi_{0}(G)\setminus\{(G)\}$ with $(\tilde{H})>(H)$. Therefore, computation of the coefficients $\operatorname{coeff}^{H}((K)\cdot(L))$ is generally impractical except for those orbit types which are maximal in $\Phi_{0}(G)$, i.e. for those orbit types $(H)\in\Phi_{0}(G)$ where $(\tilde{H})>(H)$ implies $(\tilde{H})=(G)$, in which case the recurrence formula simplifies to
$\operatorname{coeff}^{H}((K)\cdot(L))=\frac{n(H,L)|W(L)|n(H,K)|W(K)|}{|W(H)|}.$
Remark 3.1.
One of the principal difficulties when working with a product group of the form $G=O(2)\times\Gamma\times{\mathbb{Z}}_{2}$ is that the conjugacy class of the subgroup $O(2)\times\{e_{\Gamma}\}\times\{1\}\leq G$ is always a generator of the Burnside ring $A(G)$. Since any orbit type $(H)$ which is maximal in the isotropy lattice $\Phi_{0}(G;\mathscr{H}\setminus\{0\})$ must also be maximal in the sublattice $\Phi_{0}(G;\mathscr{H}_{0}\setminus\{0\})$, this implies that any non-zero function $u\in\mathscr{H}$ with an isotropy group $G_{u}\leq G$ such that $(G_{u})$ is a maximal element in $\Phi_{0}(G;\mathscr{H}\setminus\{0\})$ must be $O(2)$-invariant. Therefore, any map $u\in\mathscr{H}$ which is both non-trivial and non-constant has an isotropy group $G_{u}\leq G$ with a conjugacy class that is maximal in $\Phi_{0}(G;\mathscr{H}_{m}\setminus\{0\})$ for some positive $m$.
Motivated by Remark 3.1, the conjugacy class of a subgroup $H\leq G$ is called an orbit type of maximal kind if it is a maximal element in $\Phi_{0}(G;\mathscr{H}_{m}\setminus\{0\})$ for some $m>0$. Moreover, we denote by $\mathfrak{M}$ the set of all orbit types of maximal kind, by $\mathfrak{M}_{m}$ the set of maximal elements in $\Phi_{0}(G;\mathscr{H}_{m}\setminus\{0\})$ and by $\mathfrak{M}_{m,j}$ the set of orbit types $\mathfrak{M}_{m}\cap\Phi_{0}(G;\mathscr{H}_{m,j}\setminus\{0\})$. Since each orbit type of maximal kind $(H)\in\mathfrak{M}_{m}$ is contained in the sublattice of orbit types $\Phi_{0}(G;\mathscr{H}_{m,j}\setminus\{0\})$ for at least one $j\in\{0,1,\ldots,r\}$, we have
$\mathfrak{M}=\bigcup_{m>0}\mathfrak{M}_{m},\quad\mathfrak{M}_{m}=\bigcup_{j=0}^{r}\mathfrak{M}_{m,j},$
corresponding to the $\Gamma\times\mathbb{Z}_{2}$-isotypic decomposition of the $m$-th $O(2)$-isotypic component of $\mathscr{H}$:
$\mathscr{H}_{m}=\mathscr{H}_{m,1}\oplus\mathscr{H}_{m,2}\oplus\cdots\oplus\mathscr{H}_{m,r}.$
At this point, we can strengthen the sufficient condition for the existence of a non-trivial solution to (4) stated in Lemma 1.1 to a sufficient condition for the existence of a non-stationary solution:
Lemma 3.2.
If for some orbit type of maximal kind $(H)\in\mathfrak{M}$ one has
$\operatorname{coeff}^{H}\left(G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))\right)\neq 0,$
then there exists a non-stationary function $u\in\mathscr{H}$ satisfying (4).
3.1 The $s$-Folding Homomorphism
In this section, we introduce the concept of $s$-folding, which plays a key role in our study of orbit types of maximal kind. For each $s\in\mathbb{N}$, we define the $s$-folding homomorphism by
$\phi_{s}:O(2)\rightarrow O(2)/\mathbb{Z}_{s}\simeq O(2),\quad\phi_{s}(e^{i\theta}):=e^{is\theta},\quad\phi_{s}(\kappa e^{i\theta}):=\kappa e^{is\theta}.$
There is a natural extension of $\phi_{s}$ to the Lie group homomorphism $\psi_{s}:O(2)\times\Gamma\times\mathbb{Z}_{2}\rightarrow O(2)\times\Gamma\times\mathbb{Z}_{2}$ given by
$\psi_{s}(e^{i\theta},\gamma,\pm 1):=(\phi_{s}(e^{i\theta}),\gamma,\pm 1),\quad\psi_{s}(\kappa e^{i\theta},\gamma,\pm 1):=(\phi_{s}(\kappa e^{i\theta}),\gamma,\pm 1).$ (12)
In turn, (12) induces the Burnside ring homomorphism $\Psi_{s}:A(G)\rightarrow A(G)$ defined as follows
$\Psi_{s}(H):=({}^{s}H),\quad{}^{s}H:=\psi^{-1}_{s}(H).$ (13)
Clearly, one has the following relation between the sets $\mathfrak{M}_{m}$ and $\mathfrak{M}_{sm}$ for all $s>0$ and $m\geq 0$
$\Psi_{s}(\mathfrak{M}_{m})=\mathfrak{M}_{sm}.$
Zalman et al. hypothesize in [5]—where it serves as an essential component in the proof of their main result—that every orbit type $(H)\in\Phi_{0}(G;\mathscr{H}_{m})\setminus\{(G)\}$ satisfies the relation
$({}^{s}H)>(H)\text{ for all }s\in{\mathbb{N}}.$ (14)
In this paper, we demonstrate that the relation (14) does not hold universally. Instead, we prove that the relation between $(H)$ and $({}^{s}H)$ depends both on $(H)\in\Phi_{0}(G;\mathscr{H}_{m})$ and $s\in{\mathbb{N}}$.
4 A Classification of Generators for the Burnside ring Associated with the Group $G=O(2)\times\Gamma\times\mathbb{Z}_{2}$.
In order to proceed with our computations in the Burnside ring, we must employ the convention of amalgamated notation: a shorthand for the specification of subgroups in a product group, first considered by Balanov et al. in [8]. It is a well known consequence of Goursat’s Lemma (cf. [16]) that any closed subgroup in the product group $G=O(2)\times\Gamma\times{\mathbb{Z}}_{2}$ can be identified, up to its conjugacy class in $\Phi_{0}(G)$, with a quintuple $(K_{O},K_{\Gamma},L,\varphi_{O},\varphi_{\Gamma})$ consisting of two subgroups $K_{O}\leq O(2)$, $K_{\Gamma}\leq\Gamma\times\mathbb{Z}_{2}$, a group $L$ and a pair of epimorphisms $\varphi_{O}:K_{O}\rightarrow L$, $\varphi_{\Gamma}:K_{\Gamma}\rightarrow L$, as follows
$K_{O}{}^{\varphi_{O}}\times^{\varphi_{\Gamma}}_{L}K_{\Gamma}:=\{(x,y)\in K_{O}\times K_{\Gamma}\;:\;\varphi_{O}(x)=\varphi_{\Gamma}(y)\}.$ (15)
Let $H$ be a closed subgroup in $G$ with the amalgamated decomposition (15) and denote by $r(H)\in SO(2)$ the rotation generator in $K_{O}$. Since the order of $\varphi_{O}(r(H))\in L$ is invariant under conjugation of $r(H)$ in $K_{O}$, the following map is well-defined:
$\mathfrak{m}:\Phi_{0}(G)\rightarrow{\mathbb{N}},\quad\mathfrak{m}(H):=|\varphi_{O}(r(H))|.$ (16)
Given any orbit type $(H)\in\Phi_{0}(G)$ and a number $s_{0}\in{\mathbb{N}}$, we will demonstrate how the map (16) can be used to describe the set
$\left\{s\in{\mathbb{N}}:({}^{s}H)\leq({}^{s_{0}}H)\right\}.$ (17)
In turn, for any orbit type of maximal kind $(H)\in\mathfrak{M}$ and for any pair of generators $(K),(L)\in\Phi_{0}(G)$ we can use the set (17) to characterize the coefficients of $({}^{s}H)$ for all $s\in{\mathbb{N}}$ in the Burnside ring product
$(K)\cdot(L)\in A(G).$
Remark 4.1.
For any orbit type of maximal kind $(H)\in\mathfrak{M}$, the relation $(H)<(\tilde{H})$ implies $(\tilde{H})=(G)$ or $(\tilde{H})=({}^{s}H)$ for some $s\in{\mathbb{N}}$. Therefore, since satisfaction of the relations $(H)<(K)$ and $(H)<(L)$ is a prerequisite for the non-triviality $\operatorname{coeff}^{H}((K)\cdot(L))\neq 0$, we can restrict our focus to the coefficients of $({}^{s}H)$ for all $s\in{\mathbb{N}}$ in Burnside ring products of the form
$({}^{s_{0}}H)\cdot({}^{s_{1}}H)\in A(G),\quad s_{0},s_{1}\in{\mathbb{N}}.$
Theorem 4.2.
For any orbit type $(H)\in\Phi_{0}(G)$ and pair of numbers $s_{0},s_{1}\in{\mathbb{N}}$ with $s_{1}\leq s_{0}$, the corresponding pair of folded orbit types $({}^{s_{1}}H),({}^{s_{0}}H)\in\Phi_{0}(G)$ have the relation $({}^{s_{1}}H)\leq({}^{s_{0}}H)$ if and only if
$\mathfrak{m}(H)\mid\frac{s_{0}-s_{1}}{\gcd(s_{0},s_{1})}\text{ or}\mathfrak{m}(H)\mid\frac{s_{0}+s_{1}}{\gcd(s_{0},s_{1})}.$
Proof.
Assume that a subgroup representative $H\in(H)$ has the amalgamated decomposition $K_{O}{}^{\varphi_{O}}\times^{\varphi_{\Gamma}}_{L}K_{\Gamma}$. According to the natural ordering of $\Phi_{0}(G)$, the relation
$({}^{s_{1}}H)\leq({}^{s_{0}}H)$ (18)
is equivalent to the existence of an element $g\in G$ for which
$g\,{}^{s_{1}}Hg^{-1}\leq{}^{s_{0}}H.$ (19)
Given any $s\in{\mathbb{N}}$, put ${}^{s}K_{O}:=\phi_{s}^{-1}(K_{O})$, $\varphi^{s}_{O}:=\varphi_{O}\circ\phi_{s}:{}^{s}K_{O}\rightarrow L$ and notice that the subgroup ${}^{s}H\leq G$ has the amalgamated decomposition
$\psi_{s}^{-1}(K_{O}{}^{\varphi_{O}}\times^{\varphi_{\Gamma}}_{L}K_{\Gamma})$ $=\psi_{s}^{-1}(\{(x,y)\in K_{O}\times K_{\Gamma}\;:\;\varphi_{O}(x)=\varphi_{\Gamma}(y)\})$
=$=\{(\phi_{s}^{-1}(x),y)\in\phi_{s}^{-1}(K_{O})\times K_{\Gamma}\;:\;\varphi_{O}(x)=\varphi_{\Gamma}(y)\}$
=$=\{(x^{\prime},y)\in{}^{s}K_{O}\times K_{\Gamma}\;:\;\varphi^{s}_{O}(x^{\prime})=\varphi_{\Gamma}(y)\}$
$={}^{s}K_{O}{}^{\varphi^{s}_{O}}\times^{\varphi_{\Gamma}}_{L}K_{\Gamma}.$
On the other hand, and without loss of generality, take $g:=(a,e_{\Gamma},1)\in O(2)\times\Gamma\times\mathbb{Z}_{2}$, put $\tilde{K}_{O}:=aK_{O}a^{-1}\leq O(2)$ and define the epimorphism $\tilde{\varphi}_{O}:\tilde{K}_{O}\rightarrow L$ by
$\tilde{\varphi}_{O}(x):=\varphi_{O}(a^{-1}xa),\quad x\in\tilde{K}_{O},$
so that the subgroup $gHg^{-1}\leq G$ can be expressed, using amalgamated notation, as follows
$gK_{O}{}^{\varphi_{O}}\times^{\varphi_{\Gamma}}_{L}K_{\Gamma}g^{-1}$ $=(a,e_{\Gamma},1)\{(x,y)\in K_{O}\times K_{\Gamma}\;:\;\varphi_{O}(x)=\varphi_{\Gamma}(y)\}(a^{-1},e_{\Gamma},1)$
=$=\{(axa^{-1},y)\in aK_{O}a^{-1}\times K_{\Gamma}\;:\;\varphi_{O}(x)=\varphi_{\Gamma}(y)\}$
=$=\{(\tilde{x},y)\in\tilde{K}_{O}\times K_{\Gamma}\;:\;\tilde{\varphi}_{O}(\tilde{x})=\varphi_{\Gamma}(y)\}$
$=\tilde{K}_{O}{}^{\tilde{\varphi}_{O}}\times^{\varphi_{\Gamma}}_{L}K_{\Gamma}.$
With these notations, it becomes clear that the subgroups ${}^{s_{0}}H,{}^{s_{1}}H\leq G$ have the relation (19) for some $g=(a,\gamma,\pm 1)\in O(2)\times\Gamma\times{\mathbb{Z}}_{2}$ if and only if $\varphi^{s_{0}}_{O}(axa^{-1})=\varphi^{s_{1}}_{O}(x)$ for every $(x,y)\in{}^{\gcd(s_{0},s_{1})}K_{O}\times K_{\Gamma}$ satisfying $\varphi^{\gcd(s_{0},s_{1})}_{O}(x)=\varphi_{\Gamma}(y)$. In particular, since the $\Gamma$ component of an amalgamated subgroup is unchanged by folding, the relation (18) is equivalent to the existence of an element $a\in O(2)$ satisfying
$\varphi^{s_{1}}_{O}(axa^{-1})=\varphi^{s_{0}}_{O}(x)\text{ for all }x\in{}^{d}K_{O},\text{ where }d:=\gcd(s_{0},s_{1}).$
In turn, it is sufficient to only consider the behavior of the element $r^{\frac{1}{d}}\in{}^{d}K_{O}$ where $r$ is the rotation generator in $K_{O}$ and the two cases $a\in\{\kappa,e_{O(2)}\}$. If $a=\kappa$, then from $\kappa r\kappa=r^{-1}$, one has
$\kappa{}^{s_{1}}H\kappa\leq{}^{s_{0}}H\iff\varphi_{O}^{s_{1}}(\kappa r^{\frac{1}{d}}\kappa)=\varphi_{O}^{s_{0}}(r^{\frac{1}{d}})\iff\varphi_{O}(r)^{\frac{s_{0}+s_{1}}{d}}=e_{L}\iff\mathfrak{m}(H)\mid\frac{s_{0}+s_{1}}{d}.$
Supposing instead that $a=e_{O(2)}$, one has
${}^{s_{1}}H\leq{}^{s_{0}}H\iff\varphi_{O}^{s_{1}}(r^{\frac{1}{d}})=\varphi_{O}^{s_{0}}(r^{\frac{1}{d}})\iff\varphi_{O}(r)^{\frac{s_{0}-s_{1}}{d}}=e_{L}\iff\mathfrak{m}(H)\mid\frac{s_{0}-s_{1}}{d}.$
$\square$ For the sake of simplifying our exposition throughout the remainder of this paper, we employ the Iverson brackets which map any logical predicate $P$ to the set $\{0,1\}$ according to the rule
$[P]:=\begin{cases}1\quad&\text{ if }P\text{ is true};\\ 0\quad&\text{ otherwise},\end{cases}$ (20)
and for any set of indices $I\subset{\mathbb{N}}$, we indicate by $\mathcal{B}_{\mathfrak{m}}(I)$ the Boolean expression
[…]
$\mathfrak{m}|\frac{(x-y)}{\gcd(x,y)}\text{ or }\mathfrak{m}|\frac{(x+y)}{\gcd(x,y)}.\text{"}$ ”
Theorem 4.3.
Let $(H)\in\mathfrak{M}$ be an orbit type of maximal kind. For any pair of numbers $s_{0},s_{1}\in{\mathbb{N}}$, one has
$\operatorname{coeff}^{{}^{s}H}(({}^{s_{0}}H)\cdot({}^{s_{1}}H))=\begin{cases}|W(H)|\quad&\text{ if }s=\gcd(s_{0},s_{1})\text{ and }[\mathcal{B}_{\mathfrak{m}(H)}(\{s_{0},s_{1}\})];\\ 0\quad&\text{otherwise},\end{cases}$
equivalently, for any $s\in{\mathbb{N}}$, one has
$\operatorname{coeff}^{{}^{s}H}(({}^{s_{0}}H)\cdot({}^{s_{1}}H))=|W(H)|[s=\gcd{(s_{0},s_{1})}][\mathcal{B}_{\mathfrak{m}(H)}(\{s_{0},s_{1}\})].$
Proof.
Recall, the integer coefficient $n_{L}\in{\mathbb{Z}}$ of an orbit type $(L)$ in the Burnside ring product $({}^{s_{0}}H)\cdot({}^{s_{1}}H)$ is defined as the number of orbits of type $(L)$ contained in the $G$-space $G/{}^{s_{0}}H\times G/{}^{s_{1}}H$. In particular, $(L)\in\Phi_{0}(G)$ is an orbit type in $G/{}^{s_{0}}H\times G/{}^{s_{1}}H$ if and only if there exists a group element $g\in G$ such that
$L={}^{s_{0}}H\cap g\,{}^{s_{1}}Hg^{-1}.$
Without loss of generality, put $g=(a,e_{\Gamma},1)$ with an arbitrary $a\in O(2)$ and notice that the amalgamated decomposition
${}^{s_{0}}H\cap$ $g\,{}^{s_{1}}Hg^{-1}$
=$=\left\{(x,y)\in\left({}^{s_{0}}K_{O}\cap a{}^{s_{1}}K_{O}a^{-1}\right)\times K_{\Gamma}:\varphi_{O}^{s_{0}}(x)=\varphi_{O}^{s_{1}}(a^{-1}xa)=\varphi_{\Gamma}(y)\right\}$
$=\left({}^{s_{0}}K_{O}\cap{}^{s_{1}}\tilde{K}_{O}\right){}^{\psi}\times_{L}^{\varphi_{\Gamma}}K_{\Gamma},\quad(\text{where }{}^{s_{1}}\tilde{K}_{O}:=a{}^{s_{1}}K_{O}a^{-1})$
describes a subgroup belonging to the conjugacy class $({}^{s}H)$ if and only if there exists a group element $b\in O(2)$ with both
  1. (i)
    $\varphi_{O}(bxb^{-1})=\varphi_{O}^{s_{0}}(x)=\varphi_{O}^{s_{1}}(axa^{-1});\text{ for all }x\in K_{O},$
  2. (ii)
    $b{}^{s}K_{O}b^{-1}={}^{s_{0}}K_{O}\cap a{}^{s_{1}}K_{O}a^{-1}.$
The equivalence of condition $(\rm i)$ with the truthfulness of the Boolean expression $\mathcal{B}_{\mathfrak{m}(H)}(\{s,s_{0},s_{1}\})$ has already been demonstrated in the proof of Theorem 4.2. On the other hand, the observation
$K_{O}\in\{O(2),SO(2),D_{n},{\mathbb{Z}}_{n},n\in{\mathbb{N}}\},$
together with the facts
${\mathbb{Z}}_{s_{0}n}\cap{\mathbb{Z}}_{s_{1}n}={\mathbb{Z}}_{\gcd(s_{1},s_{0})n}\text{ and }D_{n}={\mathbb{Z}}_{n}\cup\kappa{\mathbb{Z}}_{n},$
implies that condition $(\rm ii)$ can be equated with $s=\gcd(s_{0},s_{1})$ (cf. [7], Example 3.12). For convenience, put $d:=\gcd(s_{0},s_{1})$ and notice that, in the case one has $\mathcal{B}_{\mathfrak{m}(H)}(\{d,s_{0},s_{1}\})$, the coefficient $n_{d}\in{\mathbb{Z}}$ standing next to the orbit type $({}^{d}H)$ in the Burnside ring product $({}^{s_{0}}H)\cdot({}^{s_{1}}H)\in A(G)$ can be obtained with the recurrence formula (cf. Appendix A) as follows
$n_{d}=\frac{n({}^{d}H,{}^{s_{0}}H)|W({}^{s_{0}}H)|n({}^{d}H,{}^{s_{1}}H)|W({}^{s_{1}}H)|}{|W({}^{d}H)|},$
and the result follows from the fact $|W(H)|=|W({}^{s}H)|$ for any $s\in{\mathbb{N}}$ together with the observation that $\mathcal{B}_{\mathfrak{m}(H)}(\{d,s_{0},s_{1}\})$ implies $n({}^{d}H,{}^{s_{0}}H)=n({}^{d}H,{}^{s_{1}}H)=1$. $\square$
5 A Characterization of the Orbit Types of Maximal Kind in the Product of Basic Degrees.
Since the classification put forward in the previous section fully characterizes the orbit types $(\tilde{H})$ with $(\tilde{H})>(H)$ for any orbit type of maximal kind $(H)\in\mathfrak{M}$, we are now in a position to calculate their coefficients in Burnside ring products of the form (11). Given a basic degree $\deg_{\mathcal{V}_{m,j}}\in A(G)$ corresponding to an index pair $(m,j)\in\Sigma$ with $m>0$ and an orbit type of maximal kind $(H)\in\mathfrak{M}$, the recurrence formula for the $G$-equivariant Leray-Schauder degree (cf. Appendix A) implies
$\deg_{\mathcal{V}_{m,j}}=(G)-y_{m,j}(H)+a_{m,j},$ (22)
where $a_{m,j}\in A(G)$ is such that $\operatorname{coeff}^{{}^{s}H}(a_{m,j})=0$ for all $s\in\mathbb{N}$ and the coefficient $y_{m,j}\in\mathbb{Z}$ is determined by the rule
$y_{m,j}=\begin{cases}0\quad&\text{if }\dim{\mathcal{V}_{m,j}^{H}}\text{ is even};\\ 1\quad&\text{if }\dim{\mathcal{V}_{m,j}^{H}}\text{ is odd and }|W(H)|=2;\\ 2\quad&\text{if }\dim{\mathcal{V}_{m,j}^{H}}\text{ is odd and }|W(H)|=1,\end{cases}$ (23)
or, equivalently (see discussion of the Iverson bracket notation (20)):
$y_{m,j}=x_{0}\left[2\nmid\dim{\mathcal{V}_{m,j}^{H}}\right],$
where
$x_{0}=\begin{cases}1\quad&\text{if }|W(H)|=2;\\ 2\quad&\text{if }|W(H)|=1.\end{cases}$
Lemma 5.1.
Let $(H)\in\mathfrak{M}$ be an orbit type of maximal kind. For any basic degree $\deg_{\mathcal{V}_{m,j}}$ with $(m,j)\in\Sigma$ and $m>0$ one has the implication
$\operatorname{coeff}^{H}(\deg_{\mathcal{V}_{m,j}})\neq 0\implies(H)\in\mathfrak{M}_{m,j}.$
Proof.
Since $\mathcal{V}_{m,j}^{H}=\{0\}$ for any orbit type $(H)\notin\Phi_{0}(G,\mathscr{H}_{m,j})$, the proof follows from formula (23) together with the assumption that $(H)$ is of maximal kind. $\square$ For every $s\in\mathbb{N}$, the $s$-folding homomorphism (12) induces the following relation between the basic degrees $\deg_{\mathcal{V}_{m,j}},\deg_{\mathcal{V}_{sm,j}}\in A(G)$
$\Psi_{s}(\deg_{\mathcal{V}_{m,j}})=\deg_{\mathcal{V}_{sm,j}},$ (24)
such that, for any orbit type $(H)\in\Phi_{0}(G)$, one has
$\operatorname{coeff}^{H}(\deg_{\mathcal{V}_{m,j}})=\operatorname{coeff}^{{}^{s}H}(\deg_{\mathcal{V}_{sm,j}}).$ (25)
Using the observations (22), (24), and (25), we will derive results which characterize the behavior of orbit types of maximal kind in the Burnside ring product of a finite collection of basic degrees of the form
$\left\{\deg_{\mathcal{V}_{s_{k}m,j_{k}}}\in A(G):s_{k}\in\mathbb{N}\right\}_{k=1,2,\ldots,N},\quad m>0,\;j_{k}\in\{0,1,\ldots,r\}.$ (26)
To begin with, we recall a result that the coefficient of an orbit type of maximal kind $(H)\in\mathfrak{M}$ is $2$-nilpotent with respect to the Burnside ring product of any pair of basic degrees $\deg_{\mathcal{V}_{m,j_{1}}},\deg_{\mathcal{V}_{m,j_{2}}}\in A(G)$ with $m>0$ and $j_{1},j_{2}\in\{0,1,\ldots,r\}$, provided that both $\dim\mathcal{V}_{m,j_{1}}^{H}$ and $\dim\mathcal{V}_{m,j_{2}}^{H}$ are odd. Although this fact has been demonstrated elsewhere (cf. for example [7], [8], [5], [13]), the method of its proof is instructive for the proofs of subsequent, novel results and for this reason we include our own version of the argument.
Lemma 5.2.
Let $(H)\in\mathfrak{M}$ be an orbit type of maximal kind. For any $m>0$ and $j_{1},j_{2}\in\{0,1,\ldots,r\}$, one has
$\operatorname{coeff}^{H}(\deg_{\mathcal{V}_{m,j_{1}}}\cdot\deg_{\mathcal{V}_{m,j_{2}}})=\begin{cases}0\quad&\text{if }\dim\mathcal{V}_{m,j_{1}}^{H}\text{ and }\dim\mathcal{V}_{m,j_{2}}^{H}\\ &\quad\text{ are of the same parity;}\\ -x_{0}\quad&\text{else,}\end{cases}$
or, equivalently
$\operatorname{coeff}^{H}(\deg_{\mathcal{V}_{m,j_{1}}}\cdot\deg_{\mathcal{V}_{m,j_{2}}})=-x_{0}\left[2\nmid\dim\mathcal{V}_{m,j_{1}}^{H}+\dim\mathcal{V}_{m,j_{2}}^{H}\right].$
Proof.
Consider the Burnside ring product of the relevant basic degrees
$\deg_{\mathcal{V}_{m,j_{1}}}\cdot\deg_{\mathcal{V}_{m,j_{2}}}$ $=\left((G)-y_{m,j_{1}}(H)+a_{m,j_{1}}\right)\cdot\left((G)-y_{m,j_{2}}(H)+a_{m,j_{2}}\right)$
$=(G)-(y_{m,j_{1}}+y_{m,j_{2}}-y_{m,j_{1}}y_{m,j_{2}}|W(H)|)(H)+a,$
where $a\in A(G)$ is such that $\operatorname{coeff}^{{}^{s}H}(a)=0$ for all $s\in\mathbb{N}$. Now, if $\dim\mathcal{V}_{m,j_{1}}^{H}$ and $\dim\mathcal{V}_{m,j_{2}}^{H}$ are both even, then one has $y_{m,j_{1}}=y_{m,j_{2}}=0$ and the result follows. On the other hand, if $\dim\mathcal{V}_{m,j_{1}}^{H}$ and $\dim\mathcal{V}_{m,j_{2}}^{H}$ are both odd, then one has $y_{m,j_{1}}=y_{m,j_{2}}=x_{0}$ such that
$y_{m,j_{1}}+y_{m,j_{2}}-y_{m,j_{1}}y_{m,j_{2}}|W(H)|=x_{0}(2-x_{0}|W(H)|),$
where in either of the cases: $x_{0}=2$ and $|W(H)|=1$ or $x_{0}=1$ and $|W(H)|=2$, one has $02-x_{0}|W(H)|=$ 0. Suppose instead that $\dim\mathcal{V}_{m,j_{1}}^{H}$ and $\dim\mathcal{V}_{m,j_{2}}^{H}$ are of different parties and one has the two corresponding cases, $y_{m,j_{1}}=x_{0}$ and $y_{m,j_{2}}=0$ or $y_{m,j_{1}}=0$ and $y_{m,j_{2}}=x_{0}$, both of which imply $y_{m,j_{1}}+y_{m,j_{2}}-y_{m,j_{1}}y_{m,j_{2}}|W(H)|=x_{0}$. $\square$ Since our computations in the Burnside ring often involve the product of a number of basic degrees, we consider a natural generalization of Lemma (5.2) to an orbit type of maximal kind $(H)\in\mathfrak{M}$ and the Burnside ring product of any finite collection of basic degrees $\{\deg_{\mathcal{V}_{m,j_{k}}}\}_{k=1}^{N}$ with $m>0$ and $j_{1},\ldots,j_{N}\in\{0,1,\ldots,r\}$.
Corollary 5.3.
Let $(H)\in\mathfrak{M}$ be an orbit type of maximal kind. For any finite collection of basic degrees $\{\deg_{\mathcal{V}_{m,j_{k}}}\}_{k=1}^{N}$ with $m>0$ and $j_{1},\ldots,j_{N}\in\{0,1,\ldots,r\}$, one has
$\operatorname{coeff}^{H}\left(\prod\limits_{k=1}^{N}\deg_{\mathcal{V}_{m,j_{k}}}\right)=\begin{cases}0\quad&\text{if }\dim\mathcal{V}_{m,j_{k}}^{H}\text{ is odd }\\ &\quad\text{for an even number of }j_{k}\in\{0,1,\ldots,r\};\\ -x_{0}\quad&\text{otherwise,}\end{cases}$
or, equivalently
$\operatorname{coeff}^{H}\left(\prod\limits_{k=1}^{N}\deg_{\mathcal{V}_{m,j_{k}}}\right)=-x_{0}\left[2\nmid\sum\limits_{k=1}^{N}\dim\mathcal{V}_{m,j_{k}}^{H}\right].$
The previous result only concerns the product of basic degrees associated with a fixed Fourier mode. We are also interested in characterizing the behaviour of the coefficients of an orbit type of maximal kind $(H)\in\mathfrak{M}$ and its $s$-foldings $\{({}^{s}H)\}_{s\in{\mathbb{N}}}\subset\mathfrak{M}$ in the Burnside ring product of any pair of basic degrees $\deg_{\mathcal{V}_{s_{1}m,j_{1}}},\deg_{\mathcal{V}_{s_{2}m,j_{2}}}\in A(G)$ with $m>0$, $j_{1},j_{2}\in\{0,1,\ldots,r\}$ and where $s_{1},s_{2}\in{\mathbb{N}}$ are any two numbers satisfying $s_{1}\neq s_{2}$.
Lemma 5.4.
Let $(H)\in\mathfrak{M}$ be an orbit type of maximal kind. For any pair of basic degrees $\deg_{\mathcal{V}_{s_{1}m,j_{1}}},\deg_{\mathcal{V}_{s_{2}m,j_{2}}}\in A(G)$ with $m>0$, $j_{1},j_{2}\in\{0,1,\ldots,r\}$ and $s_{1},s_{2}\in{\mathbb{N}}$ where $s_{1}\neq s_{2}$ and for any number $s_{0}\in{\mathbb{N}}$, one has
$\operatorname{coeff}^{{}^{s_{0}}H}$ $(\deg_{\mathcal{V}_{s_{1}m,j_{1}}}\cdot\deg_{\mathcal{V}_{s_{2}m,j_{2}}})=-x_{0}\left[2\nmid\dim\mathcal{V}_{m,j_{0}}^{H}\right]\left[s_{0}\in\{s_{1},s_{2}\}\right]$
$\quad+2x_{0}\left[2\nmid\dim\mathcal{V}_{m,j_{1}}^{H}\dim\mathcal{V}_{m,j_{2}}^{H}\right]\left[\operatorname{gcd}(s_{1},s_{2})=s_{0}\right][\mathcal{B}_{\mathfrak{m}(H)}(\{s_{0},s_{1},s_{2}\})],$
where
$j_{0}:=\begin{cases}j_{1}&\text{ if }s_{0}=s_{1};\\ j_{2}&\text{ if }s_{0}=s_{2}.\end{cases}$
Proof.
Consider the Burnside ring product of the relevant basic degrees
$\deg_{\mathcal{V}_{s_{1}m,j_{1}}}\cdot$ $\deg_{\mathcal{V}_{s_{2}m,j_{2}}}=\left((G)-y_{m,j_{1}}({}^{s_{1}}H)+a_{m,j_{1}}\right)\cdot\left((G)-y_{m,j_{2}}({}^{s_{2}}H)+a_{m,j_{2}}\right)$
$=(G)-y_{m,j_{1}}({}^{s_{1}}H)-y_{m,j_{2}}({}^{s_{2}}H)+y_{m,j_{1}}y_{m,j_{2}}({}^{s_{1}}H)\cdot({}^{s_{2}}H)+a$
$=(G)-y_{m,j_{1}}({}^{s_{1}}H)-y_{m,j_{2}}({}^{s_{2}}H)$ (27)
$\quad\quad\quad+y_{m,j_{1}}y_{m,j_{2}}|W(H)|({}^{d}H)[\mathcal{B}_{\mathfrak{m}(H)}(\{d,s_{1},s_{2}\})]+a$
where $d:=\operatorname{gcd}(s_{1},s_{2})$ and $a\in A(G)$ satisfies $\operatorname{coeff}^{{}^{s}H}(a)=0$ for all $s\in\mathbb{N}$. In order to demonstrate each of the boolean conditions, we consider the following cases:
  1. $(i)$
    Supposing that $s_{0}\notin\{s_{1},s_{2}\}$, if it is also the case that both $\dim\mathcal{V}_{m,j_{1}}^{H}$ and $\dim\mathcal{V}_{m,j_{2}}^{H}$ are odd and $\operatorname{gcd}(s_{1},s_{2})=s_{0}$, and $\mathcal{B}_{\mathfrak{m}(H)}(\{s_{0},s_{1},s_{2}\})$ then one has
    $\operatorname{coeff}^{{}^{s_{0}}H}(\deg_{\mathcal{V}_{s_{1}m,j_{1}}}\cdot\deg_{\mathcal{V}_{s_{2}m,j_{2}}})=x_{0}^{2}|W(H)|({}^{s_{0}}H)$
    and the result follows from the fact that $x_{0}|W(H)|=2$. On the other hand, if either of $\dim\mathcal{V}_{m,j_{1}}^{H}$ and $\dim\mathcal{V}_{m,j_{2}}^{H}$ are even or if $\operatorname{gcd}(s_{1},s_{2})\neq s_{0}$, or if $\mathcal{B}_{\mathfrak{m}(H)}(\{s_{0},s_{1},s_{2}\})$ does not hold, then the result follows immediately from the assumption that $s_{0}\notin\{s_{1},s_{2}\}$.
  2. $(ii)$
    Supposing instead that $s_{0}\in\{s_{1},s_{2}\}$, if it is also the case that $\dim\mathcal{V}_{m,j_{0}}^{H}$ is even, then one has $y_{m,j_{0}}=0$ and the result follows from (5). If one has $s_{0}\in\{s_{0},s_{1}\}$, with both $\dim\mathcal{V}_{m,j_{1}}^{H}$, $\dim\mathcal{V}_{m,j_{s_{2}}}^{H}$ odd and $s_{0}=\gcd(s_{1},s_{2})$ and $\mathcal{B}_{\mathfrak{m}(H)}(\{s_{0},s_{1},s_{2}\})$, then (5) becomes
    $\deg_{\mathcal{V}_{s_{1}m,j_{1}}}\cdot\deg_{\mathcal{V}_{s_{2}m,j_{2}}}$ $=(G)-x_{0}({}^{s_{0}}H)+2x_{0}({}^{s_{0}}H)+a^{\prime},$
    where $a^{\prime}\in A(G)$ is such that $\operatorname{coeff}^{{}^{sm}H}(a^{\prime})=0$, and the result follows. On the other hand, if either of $\dim\mathcal{V}_{s_{1}m,j_{s_{1}}}^{H},\dim\mathcal{V}_{s_{1}m,j_{s_{1}}}^{H}$ are even (in which case one has $y_{m,j_{1}}y_{m,j_{2}}=0$) or if $\operatorname{gcd}(t,l)\neq s_{0}$, or if $\mathcal{B}_{\mathfrak{m}(H)}(\{s_{0},s_{1},s_{2}\})$ does not hold, then the result follows from (5).
$\square$ What follows is the final generalization of the previous results to an orbit type of maximal kind $(H)\in\mathfrak{M}$ and its $s$-foldings $\{({}^{s}H)\}_{s\in{\mathbb{N}}}\subset\mathfrak{M}$ in the Burnside ring product of any finite collection of basic degrees $\{\deg_{\mathcal{V}_{s_{k}m,j_{s_{k}}}}\}_{k=1}^{N}$ with $m>0$, $j_{s_{1}},\ldots,j_{s_{N}}\in\{0,1,\ldots,r\}$ and where $s_{1},\ldots,s_{N}\in\mathbb{N}$ are distinct.
Theorem 5.5.
Let $(H)\in\mathfrak{M}$ be an orbit type of maximal kind. For any finite collection of basic degrees $\{\deg_{\mathcal{V}_{s_{k}m,j_{s_{k}}}}\}_{k=1}^{N}$ with $m>0$, $j_{s_{1}},\ldots,j_{s_{N}}\in\{0,1,\ldots,r\}$ and where $s_{1},\ldots,s_{N}\in\mathbb{N}$ are distinct and for any number $s_{0}\in{\mathbb{N}}$, one has
$\operatorname{coeff}^{{}^{s_{0}}H}\left(\prod\limits_{k=1}^{N}\deg_{\mathcal{V}_{s_{k}m,j_{s_{k}}}}\right)=-x_{0}\left[2\nmid\dim\mathcal{V}_{m,j_{s_{0}}}^{H}\right]\left[s_{0}\in\{s_{1},\ldots,s_{N}\}\right]$
$\quad\quad+2x_{0}\sum\limits_{\begin{subarray}{c}I\subset\{s_{1},s_{2},\ldots,s_{N}\}\\ I\neq\emptyset,\{s_{0}\}\end{subarray}}(-2)^{|I|-2}\left[\mathcal{B}_{\mathfrak{m}(H)}(I)\right]\left[s_{0}=\operatorname{gcd}(I)\right]\left[2\nmid\prod_{s\in I}\dim\mathcal{V}_{m,j_{s}}^{H}\right],$
where, if $s_{0}=s_{k}$ for any $k=1,\ldots,N$, then $j_{s_{0}}:=j_{s_{k}}$.
Proof.
Consider the Burnside ring product of the relevant basic degrees
$\prod\limits_{k=1}^{N}\deg_{\mathcal{V}_{s_{k}m,j_{s_{k}}}}$ $=\prod\limits_{k=1}^{N}(G)-y_{m,j_{s_{k}}}({}^{s_{k}}H)+b_{m,j_{s_{k}}},$ (28)
and also the related Burnside ring product
$\prod\limits_{k=1}^{N}(G)-y_{m,j_{s_{k}}}({}^{s_{k}}H).$ (29)
Notice that for any $s_{0}\in\mathbb{N}$, one has
$\operatorname{coeff}^{{}^{s_{0}}H}\left(\prod\limits_{k=1}^{N}\deg_{\mathcal{V}_{s_{k}m,j_{s_{k}}}}\right)=\operatorname{coeff}^{{}^{s_{0}}H}\left(\prod\limits_{k=1}^{N}(G)-y_{m,j_{s_{k}}}({}^{s_{k}}H)\right),$
and also that (29) has the expansion
$\prod\limits_{k=1}^{N}(G)-$ $y_{m,j_{s_{k}}}({}^{s_{k}}H)=\sum\limits_{I\subset\{s_{1},s_{2},\ldots,s_{N}\}}\prod\limits_{s\in I}-y_{m,j_{s}}({}^{s}H)$
$=\sum\limits_{I\subset\{s_{1},s_{2},\ldots,s_{N}\}}|W(H)|^{|I|-1}\left(\prod\limits_{s\in I}-y_{m,j_{s}}\right)\left[\mathcal{B}_{\mathfrak{m}(H)}(I)\right]({}^{\operatorname{gcd}(I)}H),$
where the notation $I$ is used to indicate a subset of the indices $\{s_{1},s_{2},\ldots,s_{N}\}$ and the expression $\sum_{I\subset\{s_{1},s_{2},\ldots,s_{N}\}}$ describes a summation over all such subsets, including the empty set, in which case we put $\prod_{s\in\emptyset}-y_{m,j_{s}}(\alpha_{*})({}^{s}H):=(G)$, and the full set. It follows that the coefficient of $({}^{s_{0}}H)\in\Phi_{0}(G;\mathscr{H})$ in the Burnside ring product (28), which for convenience of notation we denote by
$n_{s_{0}}:=\operatorname{coeff}^{{}^{s_{0}}H}(\prod\limits_{k=1}^{N}\deg_{\mathcal{V}_{s_{k}m,j_{s_{k}}}}),$
is specified by the formula
$n_{s_{0}}=\sum\limits_{\begin{subarray}{c}I\subset\{s_{1},s_{2},\ldots,s_{N}\}\\ I\neq\emptyset\end{subarray}}|W(H)|^{|I|-1}\left[\mathcal{B}_{\mathfrak{m}(H)}(I)\right][s_{0}=\operatorname{gcd}(I)]\prod\limits_{s\in I}-y_{m,s}.$ (30)
Now, for any non-empty subset of indices $I\subset\{s_{1},s_{2},\ldots,s_{N}\}$, the product
$\prod\limits_{s\in I}-y_{m,j_{s}}=\prod\limits_{s\in I}-x_{0}\left[2\nmid\dim\mathcal{V}_{m,j_{s}}^{H}\right],$ (31)
is determined by the rule
$\prod\limits_{s\in I}-y_{m,j_{s}}=\begin{cases}(-x_{0})^{|I|}\quad&\text{ if }\dim\mathcal{V}_{m,j_{s}}^{H}\text{ is odd for all }s\in I;\\ 0\quad&\text{ otherwise}.\end{cases}$
Therefore, (30) becomes
$\sum\limits_{\begin{subarray}{c}I\subset\{s_{1},s_{2},\ldots,s_{N}\}\\ I\neq\emptyset\end{subarray}}|W(H)|^{|I|-1}(-x_{0})^{|I|}\left[\mathcal{B}_{\mathfrak{m}(H)}(I)\right]\left[s_{0}=\operatorname{gcd}(I)\right]\left[2\nmid\prod_{s\in I}\dim\mathcal{V}_{m,j_{s}}^{H}\right]$
$=-x_{0}\left[2\nmid\dim\mathcal{V}_{m,j_{0}}^{H}\right]\left[s_{0}\in\{s_{1},\ldots,s_{N}\}\right]$
$\quad+2x_{0}\sum\limits_{\begin{subarray}{c}I\subset\{s_{1},s_{2},\ldots,s_{N}\}\\ I\neq\emptyset,\{s_{0}\}\end{subarray}}(-2)^{|I|-2}\left[\mathcal{B}_{\mathfrak{m}(H)}(I)\right]\left[s_{0}=\operatorname{gcd}(I)\right]\left[2\nmid\prod_{s\in I}\dim\mathcal{V}_{m,j_{s}}^{H}\right],$
where we have used the fact that $x_{0}|W(H)|=2$. $\square$ The following practical Corollary will prove useful in subsequent sections.
Corollary 5.6.
Let $(H)\in\mathfrak{M}$ be an orbit type of maximal kind. For any finite collection of basic degrees $\{\deg_{\mathcal{V}_{s_{k}m,j_{s_{k}}}}\}_{k=1}^{N}$ with $m>0$, $j_{s_{1}},\ldots,j_{s_{N}}\in\{0,1,\ldots,r\}$ and where $s_{1},\ldots,s_{N}\in\mathbb{N}$ are distinct and for any $s\in\{s_{1},\ldots,s_{N}\}$ with $2\nmid\dim\mathcal{V}_{m,j_{s}}^{H}$,
$\operatorname{coeff}^{{}^{s}H}$ $\left(\prod\limits_{k=1}^{N}\deg_{\mathcal{V}_{s_{k}m,j_{s_{k}}}}\right)\neq 0.$ (32)
Proof.
Indeed, notice that $-x_{0}+2x_{0}C\neq 0$ for any value of $C\in{\mathbb{N}}$. $\square$
6 On the Non-Triviality of the Degree $G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))$
Given an orbit type of maximal kind $(H)\in\mathfrak{M}$ and a positive number $s>0$, we put
$\begin{cases}\Sigma^{s}(H):=\{(m,j)\in\Sigma_{0}:2\nmid\dim\mathcal{V}_{m,j}^{{}^{s}H}\};\\ \mathfrak{n}^{s}(H):=|\Sigma^{s}(H)|,\end{cases}$
and
$S(H):=\{s\in{\mathbb{N}}:\mathfrak{n}^{s}(H)\neq 0\}.$
Remark 6.1.
If $(H)\in\mathfrak{M}_{m_{0}}$ for some $m_{0}>0$, then $(m,j)\in\Sigma^{s}(H)$ only if $m=m_{0}s$.
We are now in a position to present our main result.
Proposition 6.2.
Let $(H)\in\mathfrak{M}$ be an orbit type of maximal kind. For any $s_{0}\in{\mathbb{N}}$, one has
$\operatorname{coeff}^{{}^{s_{0}}H}$ $\left(G\text{\rm-deg}\left(\mathscr{A},B(\mathscr{H})\right)\right)=-x_{0}\left[2\nmid\mathfrak{n}^{s_{0}}(H)\right]\left[s_{0}\in S(H)\right]$
$\quad+2x_{0}\sum\limits_{\begin{subarray}{c}I\subset S(H)\\ I\neq\emptyset,\{s_{0}\}\end{subarray}}(-2)^{|I|-2}\left[\mathcal{B}_{\mathfrak{m}(H)}(I)\right]\left[s_{0}=\operatorname{gcd}(I)\right]\left[2\nmid\prod_{s\in I}\mathfrak{n}^{s}(H)\right].$
Proof.
For convenience of notation, put
$\Sigma^{0}(H):=\{(m,j)\in\Sigma_{0}:2\mid\dim\mathcal{V}_{m,j}^{{}^{s}H}\text{for all }s>0\},$
and define
$\rho^{s}_{G}(H):=\prod_{(m,j)\in\Sigma^{s}(H)}\deg_{\mathcal{V}_{m,j}}.$
Enumerating the set $S(H)=\{s_{1},s_{2},\ldots,s_{N}\}$, one clearly has
$\Sigma_{0}\setminus\Sigma^{0}(H)=\bigcup_{k=1}^{N}\Sigma^{s_{k}}(H),$
such that the degree computation (11) becomes
$G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))=\rho^{0}_{G}(H)\cdot\prod_{k=1}^{N}\rho^{s_{k}}_{G}(H).$
From Lemma 5.2 and Corollary 5.3, one has
$\rho_{G}^{s}(H)=\begin{cases}(G)+b_{0}&\text{ if }s=0;\\ (G)-y_{s}({}^{s}H)+b_{s}&\text{ if }s>0,\end{cases}$
where $b_{0},b_{s}$ satisfy $\operatorname{coeff}^{{}^{s}H}(b_{0})=\operatorname{coeff}^{{}^{s}H}(b_{s})=0$ for all $s\in{\mathbb{N}}$ and
$y_{s}=\begin{cases}0&\text{ if }\mathfrak{n}^{s}(H)\text{ is even};\\ x_{0}&\text{ if }\mathfrak{n}^{s}(H)\text{ is odd}.\end{cases}$
Assume, without loss of generality, that $(H)\in\mathfrak{M}_{m_{0}}$ and notice that each of the elements $\rho_{G}^{s_{k}}(H)$ has the form of a basic degree corresponding to an index $(s_{k}m_{0},j_{s_{k}})\in\Sigma$ satisfying
$2\mid\dim\mathcal{V}_{m_{0},j_{s_{k}}}^{H}+\mathfrak{n}^{s_{k}}(H),$
i.e., such that
$\operatorname{coeff}^{{}^{s_{k}}H}\left(\deg_{\mathcal{V}_{s_{k}m_{0},j_{s_{k}}}}\right)=\operatorname{coeff}^{{}^{s_{k}}H}\left(\rho_{G}^{s_{k}}(H)\right).$
Therefore, the coefficient of any $s_{0}$-folding $({}^{s_{0}}H)\in\mathfrak{M}$ in the degree (11) can be computed as follows
$\operatorname{coeff}^{{}^{s_{0}}H}(G\text{\rm-deg}(\mathscr{A},B(\mathscr{H})))=\operatorname{coeff}^{{}^{s_{0}}H}\left(\prod_{k=1}^{N}\deg_{\mathcal{V}_{s_{k}m_{0},j_{s_{k}}}}\right),$
and the result follows from Theorem 5.5. $\square$ Just as Corollary 5.6 naturally follows from Theorem 5.5, the following Corollary is a direct implication of Proposition 6.2.
Corollary 6.3.
Let $(H)\in\mathfrak{M}$ be an orbit type of maximal kind. For any $s\in{\mathbb{N}}$ with $2\nmid n^{s}(H)$, one has
$\operatorname{coeff}^{{}^{s}H}\left(G\text{\rm-deg}\left(\mathscr{A},B(\mathscr{H})\right)\right)\neq 0.$
In turn, combining Lemma 1.1 with Corollary 6.3 leads immediately to Theorem 1.2.
7 Motivating Example: An Arrangement of $N$ Coupled Pendula with Dihedral Symmetries and Subject to Nonlinear Forcing
Figure 1: Cycle of $N$ oscillating pendula with $\Gamma=D_{N}$ symmetries
We can provide a more definitive prediction of the symmetries expressed by non-stationary solutions to (1) than that established in Theorem 1.2 by prescribing an exact coupling relation between the pendula in our configuration. Specifically, let $\Gamma$ be the dihedral group of order $2N$ such that our full symmetry group becomes
$G:=O(2)\times D_{N}\times{\mathbb{Z}}_{2}.$
Our model for this configuration is the system
$\begin{cases}\ddot{u}(t)=|u|^{q}u-(L+\operatorname{Id})u,\quad t\in\mathbb{R},\;u(t)\in V\\ u(t)=u(t+p),\;\dot{u}(t)=\dot{u}(t+p),\end{cases}$ (33)
where $|u|^{q}u:=(|u|^{q}u_{1},|u|^{q}u_{2},\ldots,|u|^{q}u_{N})$ for any even $q>1$ and $L:V\rightarrow V$ is the graph Laplacian matrix for an undirected graph $\mathbb{G}$ invariant under the permutation action of $D_{N}$, with $N$ vertices representing a collection of the same number of oscillating pendula and with edges representing the coupling relations between vertices. Since the natural permutation representation of $\Gamma$ on $V$
$\rho_{V}:\Gamma\rightarrow GL(V),\quad\rho_{V}(\sigma)(u_{1},u_{2},\ldots,u_{N}):=(u_{\sigma(1)},u_{\sigma(2)},\ldots,u_{\sigma(N)}),$ (34)
is induced by the permutation actions of the rotation and reflection generators $\gamma,\kappa\in\Gamma$ on the indices $i\in\{1,\ldots,N\}$
$\gamma(i):=i+1\pmod{N}\text{ and }\kappa(i):=N-i\pmod{N},$
its character can be determined according to the rule
$\chi_{V}(\gamma)=0\;\text{ and }\;\chi_{V}(\kappa)=\begin{cases}1&\text{ if }N\text{ is odd};\\ 0&\text{ if }N\text{ is even}.\end{cases}$
The number and character of the irreducible $\Gamma$ representations also depend on the dihedral order $N$: there are always the trivial representation $\mathcal{V}_{0}\simeq\mathbb{R}$, on which $\Gamma$ acts trivially, $\lfloor\frac{N+1}{2}\rfloor-1$ geometric representations $\mathcal{V}_{j}$, each with an action induced by the corresponding matrix representation $\rho_{j}:\Gamma\rightarrow GL({\mathbb{C}})$
$\rho_{j}(\kappa):=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix},\quad\rho_{j}(\gamma):=\begin{pmatrix}\cos(\frac{2j\pi}{N})&-\sin(\frac{2j\pi}{N})\\ \sin(\frac{2j\pi}{N})&\cos(\frac{2j\pi}{N})\end{pmatrix},\quad 1\leq j<\lfloor\frac{N+1}{2}\rfloor,$
the sign representation $\mathcal{V}_{*}\simeq\mathbb{R}$, with the action
$\sigma\cdot v:=\text{\rm sign\,}(\sigma)v,\quad v\in\mathcal{V}_{*},$
and, in the case that $N$ is even, two additional irreducible one-dimensional representations $\mathcal{V}_{\lfloor\frac{N+1}{2}\rfloor},\mathcal{V}_{**}\simeq\mathbb{R}$ equipped, respectively, with the actions
$\sigma\cdot v:=-\text{\rm sign\,}(\sigma)v,\quad v\in\mathcal{V}_{\lfloor\frac{N+1}{2}\rfloor},$
and
$\rho_{**}(\kappa)=\rho_{**}(\gamma)=-1.$
Comparing characters for the irreducible representations of $\Gamma$ with the character of $V$
conjugacy classes $e_{\Gamma}$ $\kappa$ $\gamma$
$\chi_{0}$ $1$ $1$ $1$
$\chi_{1}$ $2$ 00 $\cos(\frac{2\pi}{N})+\cos(\frac{2\pi}{N})$
$\vdots$ $\vdots$ $\vdots$ $\vdots$
$\chi_{j}$ $2$ 00 $\cos(\frac{2j\pi}{N})+\cos(\frac{2j\pi}{N})$
$\vdots$ $\vdots$ $\vdots$ $\vdots$
$\chi_{\lfloor\frac{N+1}{2}\rfloor}$ $1$ $-1$ $1$
$\chi_{*}$ $1$ $1$ $-1$
$\chi_{**}$ $1$ $-1$ $-1$
$\chi_{V}$ $N$ $\frac{1}{2}(1-(-1)^{N})$ 00
Table 1: Character Table for $D_{N}$.
one obtains the relation
$\chi_{V}=\begin{cases}\chi_{0}+\chi_{1}+\cdots+\chi_{\frac{N+1}{2}-1}&\text{if }2\nmid N;\\ \chi_{0}+\chi_{1}+\cdots+\chi_{\frac{N}{2}-1}+\chi_{\frac{N}{2}}&\text{if }2\mid N,\end{cases}$ (35)
implying that $V$ has the $\Gamma$-isotypic decomposition
$V=\begin{cases}V_{0}\oplus V_{1}\oplus\cdots\oplus V_{\frac{N+1}{2}-1}&\text{if }2\nmid N;\\ V_{0}\oplus V_{1}\oplus\cdots\oplus V_{\frac{N}{2}-1}\oplus V_{\frac{N}{2}}&\text{if }2\mid N.\end{cases}$ (36)
For the sake of generality, we adopt the following notation for the set of $\Gamma$-isotypic indices relevant to the $\Gamma$-isotypic decomposition of $V$
$\mathfrak{J}(N):=\begin{cases}\{0,1,\ldots,\frac{N+1}{2}-1\}&\text{if }2\nmid N;\\ \{0,1,\ldots,\frac{N}{2}-1,\frac{N}{2}\}&\text{if }2\mid N.\end{cases}$
The graph Laplacian associated with the undirected graph $\mathbb{G}$ with $N$ vertices invariant under the permutation action of $D_{N}$ has the form
$L=\left(\begin{array}[c]{ccccccc}-2&1&0&\dots&0&0&1\\ 1&-2&1&\dots&0&0&0\\ 0&1&-2&\dots&0&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots&\vdots\\ 0&0&0&\dots&-2&1&0\\ 0&0&0&\dots&1&-2&1\\ 1&0&0&\dots&0&1&-2\end{array}\right).$
One can verify that such a matrix has the eigenvectors
$v_{j}:=(1,\gamma^{j},\gamma^{2j},\ldots,\gamma^{(N-1)j})^{T},\quad j\in\{0,1,\ldots,\lfloor\frac{N+1}{2}\rfloor-1\},\;\gamma:=e^{\frac{2\pi i}{N}}$
corresponding to the eigenvalues
$-z_{j}:=-2+\gamma^{j}+\gamma^{-j}=-4\sin^{2}\left(\frac{\pi j}{N}\right).$
And, in the case that $N$ is even, one will find that there is also the eigenpair
$v_{\frac{N}{2}}:=(1,-1,1\ldots,-1)^{T},\quad z_{\frac{N}{2}}=-4.$
Since each isotypic component of $V$ has simple isotypic multiplicity, i.e. since
$m_{j}=1,\quad j\in\mathfrak{J}(N),$
the eigenvalues of $\mathscr{A}$ (cf. Definition (7)) have the form
$\mu_{m,j}=\frac{m^{2}-\beta^{2}(z_{j}+1)}{1+m^{2}}.$
Therefore, the condition $\mu_{m,j}<0$ is equivalent to the condition
$14\sin^{2}\left(\frac{\pi j}{N}\right)>\frac{m^{2}}{\beta^{2}}+$ 1
and Theorem 1.2 becomes:
Theorem 7.1.
Take $m>0$ and let $(H)$ be a maximal element in the isotropy lattice $\Phi_{0}(G;\mathscr{H}_{m}\setminus\{0\})$. For the symmetry group $\Gamma=D_{N}$ and under the assumptions ($A_{0}$)($A_{4}$), if the relations
$4\sin^{2}\left(\frac{\pi j}{N}\right)>\frac{m^{2}}{\beta^{2}}+1\text{ and }2\nmid\dim\mathcal{V}_{m,j}^{H},$
are satisfied for an odd number of $D_{N}$ isotypic indices $j\in\mathfrak{J}(N)$, then there exists a non-stationary solution $u\in\mathscr{H}\setminus\{0\}$ to the system (33) with an isotropy subgroup $G_{u}\leq G$ satisfying $(G_{u})\geq(H)$.
7.1 The Special Case of $\Gamma=D_{8}$
While the framework described in Section 7 applies to any number of pendula, we can explicitly identify spatio-temporal symmetries of the non-stationary solutions to (33) by choosing any particular dihedral order. Let’s examine the special case of $8$ pendula, such that our full symmetry group becomes
$G:=O(2)\times D_{8}\times{\mathbb{Z}}_{2}.$
Since the non-triviality of the coefficient standing next to an orbit type of maximal kind $(H)\in\mathfrak{M}_{m}$ in the basic degree $\deg_{\mathcal{V}_{m,j}}\in A(G)$ is equivalent to the odd-dimensionality of the associated $H$-fixed point space $\mathcal{V}_{m,j}^{H}$, the full power of Theorem 7.1 can be demonstrated using a complete list of maximal elements in the isotropy lattice $\Phi_{0}(G;\mathscr{H}_{m}\setminus\{0\})$ for some fixed $m>0$ together with computations of the relevant basic degrees $\deg_{\mathcal{V}_{m,j}}$ for $j\in\{0,1,2,3,4\}$. This can be achieved for $m=1$ (and then extended to any $m\in{\mathbb{N}}$ via $s$-folding, cf. Section 3.1) by the following GAP code:
1# A G.A.P. Program for the computation of Maximal Orbit Types and Basic Degrees associated with the G-isotypic decomposition
2# V = V_0 $\times$ V_1 $\times$ V_2 $\times$ V_3 $\times$ V_4
3LoadPackage ("EquiDeg");
4# create groups O(2)xD8xZ2
5o2:=OrthogonalGroupOverReal(2);
6d8:=pDihedralGroup(8);
7z2:=pCyclicGroup(2);
8# generate D8xZ2
9g1:=DirectProduct(d8,z2);
10# set names for subgroup conjugacy classes in D8xZ2
11SetCCSsAbbrv(g1,["Z1","Z2","D1t","D1z","D1","Z1m","Z1p","D1zt","D1pt","D1p","D2","Z4","D2t","D2zt","D2d","Z4d","D2dt","D2z","Z2p","Z4p","D4dt","D2p","D4","D2pt","D4z","D4zt","Z8","Z8d","D4t","D4d","D4p","Z8p","D8","D4pt","D8d","D8z","D8dt","D8p"]);
12# generate O(2)xD8xZ2
13G:=DirectProduct(o2,g1);
14ccs:=ConjugacyClassesSubgroups(G);
15# find Maximal Orbit Types and Basic Degrees in first O(2)-isotypic component (these characterize the maximal orbit types in all O(2)-isotypic components via s-folding)
16irrs := Irr(G);
17# The D_8-isotypic components 0,1,2,3,4 are indexed in GAP as 1,9,11,13,7.
18#Looping through these indices we obtain the desired data as follows:
19for i in [1,9,11,13,7] do
[…]
22 max_orbtyps := MaximalOrbitTypes( irrs[1,i] );
23 basic_deg := BasicDegree( irrs[1,i] );
24 View(basic_deg);
[…]
26 View(max_orbtyps);
27od;
[…]
In the GAP package EquiDeg, developed for equivariant degree computations in the Burnside ring (cf. [26]), the conjugacy class of an amalgamated subgroup (cf. Section 4)
$H=K_{O}{}^{\varphi_{O}}\times^{\varphi_{\Gamma}}_{L}K_{\Gamma},$
is identified by the quadruple $(K_{1},K_{2},Z_{1},Z_{2})$ where $Z_{1}:=\text{\rm Ker\,}\varphi_{1}$ and $Z_{2}:=\text{\rm Ker\,}\varphi_{2}$ with the notation
$(H)=:(K_{1}{}^{Z_{1}}\times^{Z_{2}}K_{2}).$
Using this modified amalgamated notation, we can express the output of the above program, i.e. the maximal elements in the isotropy lattices $\Phi_{0}(G;\mathscr{H}_{m,j}\setminus\{0\})$ for $j\in\{0,1,2,3,4\}$ and the corresponding basic degrees as follows:
$\mathfrak{M}_{m,0}$ =$=\left\{(D_{m}\times D_{8}^{p})\right\};$
$\mathfrak{M}_{m,1}$ =$=\left\{(D_{4m}{}^{{\mathbb{Z}}_{m}}\times^{{\mathbb{Z}}_{4}^{d}}D_{8}^{p}),\;(D_{2m}{}^{D_{m}}\times^{\tilde{D}_{4}^{d}}\tilde{D}_{4}^{p}),\;(D_{2m}{}^{D_{m}}\times^{D_{4}^{d}}D_{4}^{p})\right\};$
$\mathfrak{M}_{m,2}$ =$=\left\{(D_{2m}{}^{D_{m}}\times^{D_{2}^{d}}D_{2}^{p}),\;(D_{2m}{}^{D_{m}}\times^{\tilde{D}_{2}^{d}}\tilde{D}_{2}^{p}),\;(D_{8m}{}^{{\mathbb{Z}}_{m}}\times^{{\mathbb{Z}}_{2}^{-}}D_{8}^{p})\right\};$
$\mathfrak{M}_{m,3}$ =$=\left\{(D_{2m}{}^{D_{m}}\times^{D_{1}^{p}}D_{2}^{p}),\;(D_{2m}{}^{D_{m}}\times^{\tilde{D}_{1}^{p}}\tilde{D}_{2}^{p}),\;(D_{8m}{}^{{\mathbb{Z}}_{m}}\times^{{\mathbb{Z}}_{1}^{p}}D_{8}^{p})\right\};$
$\mathfrak{M}_{m,4}$ =$=\left\{(D_{2m}{}^{D_{m}}\times^{\tilde{D}_{4}^{p}}D_{8}^{p})\right\},$
and
$\deg_{\mathcal{V}_{m,0}}$ $=(G)-(D_{m}\times D_{8}^{p});$
$\deg_{\mathcal{V}_{1,1}}$ $=(G)+2(D_{2m}{}^{{\mathbb{Z}}_{m}}\times^{{\mathbb{Z}}_{4}^{d}}\tilde{D}_{4}^{p})+2(D_{2m}{}^{{\mathbb{Z}}_{m}}\times^{{\mathbb{Z}}_{4}^{d}}D_{4}^{p})+(D_{2m}{}^{D_{m}}\times^{{\mathbb{Z}}_{4}^{d}}{\mathbb{Z}}_{4}^{p})$
$\quad\quad-2(D_{4m}{}^{{\mathbb{Z}}_{m}}\times^{{\mathbb{Z}}_{4}^{d}}D_{8}^{p})-(D_{2m}{}^{D_{m}}\times^{\tilde{D}_{4}^{d}}\tilde{D}_{4}^{p})-(D_{2m}{}^{D_{m}}\times^{D_{4}^{d}}D_{4}^{p});$
$\deg_{\mathcal{V}_{1,2}}$ $=(G)+2(D_{2m}{}^{{\mathbb{Z}}_{m}}\times^{{\mathbb{Z}}_{2}^{-}}D_{2}^{p})+2(D_{2m}{}^{{\mathbb{Z}}_{m}}\times^{{\mathbb{Z}}_{2}^{-}}\tilde{D}_{2}^{p})+(D_{2m}{}^{D_{m}}\times^{{\mathbb{Z}}_{2}^{-}}{\mathbb{Z}}_{2}^{p})$
$\quad\quad-(D_{2m}{}^{D_{m}}\times^{D_{2}^{d}}D_{2}^{p})-(D_{2m}{}^{D_{m}}\times^{\tilde{D}_{2}^{d}}\tilde{D}_{2}^{p})-2(D_{8m}{}^{{\mathbb{Z}}_{m}}\times D_{8}^{p});$
$\deg_{\mathcal{V}_{1,3}}$ $=(G)+2(D_{2m}{}^{{\mathbb{Z}}_{m}}\times^{{\mathbb{Z}}_{1}^{p}}D_{2}^{p})+2(D_{2m}{}^{{\mathbb{Z}}_{m}}\times^{{\mathbb{Z}}_{1}^{p}}\tilde{D}_{2}^{p})+(D_{2m}{}^{D_{m}}\times^{{\mathbb{Z}}_{1}^{p}}{\mathbb{Z}}_{2}^{p})$
$\quad\quad-(D_{2m}{}^{D_{m}}\times^{D_{1}^{p}}D_{2}^{p})-(D_{2m}{}^{D_{m}}\times^{\tilde{D}_{1}^{d}}\tilde{D}_{2}^{p})-2(D_{8m}{}^{{\mathbb{Z}}_{m}}\times^{{\mathbb{Z}}_{1}^{p}}D_{8}^{p});$
$\deg_{\mathcal{V}_{m,0}}$ $=(G)-(D_{2m}{}^{D_{m}}\times^{\tilde{D}_{4}^{p}}D_{8}^{p}).$
In addition to the notation defined above, we have adopted the following ancillary shorthand for identification of subgroups in $D_{8}$ generated by the rotation $\gamma:=e^{\frac{2\pi i}{8}}$ and the relection $\kappa$:
${\mathbb{Z}}_{1}:=\{1\},\quad{\mathbb{Z}}_{2}:=\{1,\gamma^{4}\},\quad{\mathbb{Z}}_{4}:=\{1,\gamma^{2},\gamma^{4},\gamma^{6}\},$
$D_{1}:=\{1,\kappa\},\quad\tilde{D}_{1}:=\{1,\kappa\gamma\},\quad D_{2}:=\{1,\gamma^{4},\kappa,\kappa\gamma^{4}\},$
$\tilde{D}_{2}:=\{1,\gamma^{4},\kappa\gamma,\kappa\gamma^{5}\},\quad D_{4}:=\{1,\gamma^{2},\gamma^{4},\gamma^{6},\kappa,\kappa\gamma^{2},\kappa\gamma^{4},\kappa\gamma^{6}\},$
$\tilde{D}_{4}:=\{1,\gamma^{2},\gamma^{4},\gamma^{6},\kappa\gamma,\kappa\gamma^{3},\kappa\gamma^{5},\kappa\gamma^{7}\},$
$D_{8}:=\{1,\gamma,\gamma^{2},\gamma^{3},\gamma^{4},\gamma^{5},\gamma^{6},\gamma^{7},\kappa,\kappa\gamma,\kappa\gamma^{2},\kappa\gamma^{3},\kappa\gamma^{4},\kappa\gamma^{5},\kappa\gamma^{6},\kappa\gamma^{7}\}$
Likewise, the reader can consult the following list of shorthand to identify the subgroups in $D_{8}\times\mathbb{Z}_{2}$:
${\mathbb{Z}}_{1}^{p}=\{(1,1),(1,-1)\},\quad{\mathbb{Z}}_{2}^{-}=\{(1,1),(\gamma^{4},-1)\},\quad{\mathbb{Z}}_{4}^{d}:=\{(1,1),(\gamma^{2},-1),(\gamma^{4},1),(\gamma^{6},-1)\},$
$D_{1}^{p}:=\{(1,1),(\kappa,1),(1,-1),(\kappa,-1)\},\quad\tilde{D}_{1}^{p}:=\{(1,1),(\kappa\gamma,1),(1,-1),(\kappa\gamma,-1)\},$
$D_{2}^{d}:=\{(1,1),(\gamma^{4},-1),(\kappa,1),(\kappa\gamma^{4},-1)\},\quad\tilde{D}_{2}^{d}:=\{(1,1),(\gamma^{4},-1),(\kappa\gamma,1),(\kappa\gamma^{5},-1)\},$
$D_{2}^{p}:=\{(1,1),(\gamma^{4},1),(\kappa,1),(\kappa\gamma^{4},1),(1,-1),(\gamma^{4},-1),(\kappa,-1),(\kappa\gamma^{4},-1)\},$
$\tilde{D}_{2}^{p}:=\{(1,1),(\gamma^{4},1),(\kappa\gamma,1),(\kappa\gamma^{5},1),(1,-1),(\gamma^{4},-1),(\kappa\gamma,-1),(\kappa\gamma^{5},-1)\},$
$D_{4}^{d}:=\{(1,1),(\gamma^{2},-1),(\gamma^{4},1),(\gamma^{6},-1),(\kappa,1),(\kappa\gamma^{2},-1),(\kappa\gamma^{4},1),(\kappa\gamma^{6},-1)\}$
$\tilde{D}_{4}^{d}:=\{(1,1),(\gamma^{2},-1),(\gamma^{4},1),(\gamma^{6},-1),(\kappa\gamma,1),(\kappa\gamma^{3},-1),(\kappa\gamma^{5},1),(\kappa\gamma^{7},-1)\}$
$\tilde{D}_{4}^{p}:=\{(1,1),(\gamma^{2},1),(\gamma^{4},1),(\gamma^{6},1),(\kappa\gamma,1),(\kappa\gamma^{3},1),(\kappa\gamma^{5},1),(\kappa\gamma^{7},1),$
$\quad\quad\quad\quad(1,-1),(\gamma^{2},-1),(\gamma^{4},-1),(\gamma^{6},-1),(\kappa\gamma,-1),(\kappa\gamma^{3},-1),(\kappa\gamma^{5},-1),(\kappa\gamma^{7},-1),\}$
$D_{8}^{p}:=\{(1,1),(\gamma,1),(\gamma^{2},1),(\gamma^{3},1),(\gamma^{4},1),(\gamma^{5},1),(\gamma^{6},1),(\gamma^{7},1),(\kappa,1),(\kappa\gamma,1),(\kappa\gamma^{2},1),$
$\quad\quad\quad\quad(\kappa\gamma^{3},1),(\kappa\gamma^{4},1),(\kappa\gamma^{5},1),(\kappa\gamma^{6},1),(\kappa\gamma^{7},1),(1,-1),(\gamma,-1),(\gamma^{2},-1),(\gamma^{3},-1),$
$\quad\quad\quad\quad(\gamma^{6},-1),(\gamma^{7},-1),(\kappa,-1),(\kappa\gamma,-1),(\kappa\gamma^{2},-1),(\kappa\gamma^{3},-1),(\kappa\gamma^{4},-1),(\kappa\gamma^{5},-1),$
$\quad\quad\quad\quad(\kappa\gamma^{6},-1),(\gamma^{4},-1),(\gamma^{5},-1),(\kappa\gamma^{7},-1)\}$
At this point, Proposition 1.4 follows directly from the observations $(\rm i)$ that the fixed point set $\mathcal{V}_{m,j}^{H}$ is odd dimensional for any $(H)\in\mathfrak{M}_{m,j}$ and $(\rm ii)$ that the sets $\mathfrak{M}_{m,j}$, $m>0$, $j\in\{0,1,2,3,4\}$ are pairwise disjoint.
Appendix A The $G$-Equivariant Leray-Schauder Degree
Given an isometric Banach $G$-representation $\mathscr{H}$ of functions taking values in an orthogonal $\Gamma$-representation $V$, the $G$-equivariant Leray-Schauder degree is a topological tool used to study the solution sets associated with equations of the form
$\mathscr{F}(u)=0,\quad u\in\mathscr{H},$ (37)
where $\mathscr{F}:\mathscr{H}\rightarrow\mathscr{H}$ is an operator satisfying the following conditions:
  1. ($B_{1}$)
    $\mathscr{F}$ is a $G$-equivariant completely continuous field;
  2. ($B_{2}$)
    there exists a sufficiently large $R>0$ such that $\mathscr{F}$ is $B_{R}(0)$-admissibly $G$-homotopic to the identity operator $\operatorname{Id}:\mathscr{H}\rightarrow\mathscr{H}$;
  3. ($B_{3}$)
    $D\mathscr{F}(0):\mathscr{H}\rightarrow\mathscr{H}$ exists and the operator $\operatorname{Id}-D\mathscr{F}(0):\mathscr{H}\rightarrow\mathscr{H}$ is a $G$-equivariant compact field;
  4. ($B_{4}$)
    if $D\mathscr{F}(0):\mathscr{H}\rightarrow\mathscr{H}$ is an isomorphism, there exists a sufficiently small $\epsilon>0$ such that $\mathscr{F}$ is $B_{\epsilon}(0)$-admissibly $G$-homotopic to $D\mathscr{F}(0)$.
Condition ($B_{1}$) allows the problem of the existence of solutions for equation (37) to be reformulated as a question concerning the non-triviality of the degree $G\text{\rm-deg}(\mathscr{F},B(\mathscr{H}))$. In turn, Conditions ($B_{2}$)-($B_{4}$) reduce calculation of $G\text{\rm-deg}(\mathscr{F},B(\mathscr{H}))$ to a Burnside ring product involving a finite number of computationally simpler $G$-basic degrees. Equivariant notation. Let $G$ be a compact Lie group. For any subgroup $H\leq G$, we denote by $(H)$ its conjugacy class, by $N(H)$ its normalizer by $W(H):=N(H)/H$ its Weyl group in $G$. The set of all subgroup conjugacy classes in $G$ is denoted by $\Phi(G):=\{(H):H\leq G\}$ and has a natural partial order defined as follows
$(H)\leq(K)\iff\exists_{g\in G}\;\;gHg^{-1}\leq K.$
As is possible with any partially ordered set, we extend the natural order over $\Phi(G)$ to a total order, which we indicate by $<$ to differentiate the two relations. Moreover, we put $\Phi_{0}(G):=\{(H)\in\Phi(G):\text{$W(H)$ is finite}\}$ and, for any $(H),(K)\in\Phi_{0}(G)$, we denote by $n(H,K)$ the number of subgroups $\tilde{K}\leq G$ with $\tilde{K}\in(K)$ and $H\leq\tilde{K}$. Given a $G$-space $X$ with an element $x\in X$, we denote by $G_{x}:=\{g\in G:gx=x\}$ the isotropy group of $x$ and we call $(G_{x})\in\Phi(G)$ the orbit type of $x\in X$. Put $\Phi(G,X):=\{(H)\in\Phi_{0}(G):(H)=(G_{x})\;\text{for some $x\in X$}\}$ and $\Phi_{0}(G,X):=\Phi(G,X)\cap\Phi_{0}(G)$. For a subgroup $H\leq G$, the subspace $X^{H}:=\{x\in X:G_{x}\geq H\}$ is called the $H$-fixed-point subspace of $X$. If $Y$ is another $G$-space, then a continuous map $f:X\to Y$ is said to be $G$-equivariant if $f(gx)=gf(x)$ for each $x\in X$ and $g\in G$. The Burnside ring. The free $\mathbb{Z}$-module $A(G):=\mathbb{Z}[\Phi_{0}(G)]$ has a natural ring structure when equipped with the multiplicative operation defined, for any pair of generators $(H),(K)\in\Phi_{0}(G)$, as follows
$(H)\cdot(K):=\sum\limits_{(L)\in\Phi_{0}(G)}n_{L}(L),$
and where the coefficients $n_{L}\in\mathbb{Z}$ are given by the recurrence formula
$n_{L}:=\frac{n(L,H)|W(H)|n(L,K)|W(K)|-\sum_{(\tilde{L})>(L)}n_{\tilde{L}}n(L,\tilde{L})|W(\tilde{L})|}{|W(L)|}.$
Any Burnside ring element $a\in A(G)$ can be expressed as a formal sum over some finite number of generator elements
$a=n_{1}(H_{1})+n_{2}(H_{2})+\cdots+n_{N}(H_{N}),$
and we use the notation
$\operatorname{coeff}^{H}(a)=n_{H},$
to specify the integer coefficient standing next to the generator element $(H)\in\Phi_{0}(G)$. An Axiomatic Construction of the $G$-equivariant Leray-Schauder-Equivariant degree. Let $\mathscr{E}$ be any isometric Banach $G$-representation. A map $f:\mathscr{E}\rightarrow\mathscr{E}$ is said to be a completely continuous $G$-equivariant field if it can be expressed in the form
$f(x)=x-F(x),$
for some compact $G$-equivariant map $F:\mathscr{E}\rightarrow\mathscr{E}$. Moreover, given an open bounded $G$-invariant set $\Omega\subset\mathscr{E}$, a completely continuous $G$-equivariant field is said to be $\Omega$-admissible and the pair $(f,\Omega)$ is called an admissible $G$-pair if $f(x)\neq 0$ for all $x\in\partial\Omega$. We denote by $\mathscr{M}^{G}(\mathscr{E})$ the set of all admissible $G$-pairs in $\mathscr{E}$ and by $\mathscr{M}^{G}$ the set of all admissible $G$-pairs defined by taking a union over all isometric Banach $G$-representations as follows
$\mathscr{M}^{G}:=\bigcup_{\mathscr{E}}\mathscr{M}^{G}(\mathscr{E}).$
The $G$-equivariant Leray-Schauder degree is defined as the unique map $G\text{\rm-deg}:\mathscr{M}^{G}\rightarrow A(G)$ that assigns to every admissible $G$-pair $(f,\Omega)$ the Burnside ring element
$G\text{\rm-deg}(f,\Omega)=\sum_{(H)\in\Phi_{0}(G)}{n_{H}(H)},$ (38)
satisfying the four degree axioms
  • (Existence) If $n_{H}\neq 0$ for some $(H)\in\Phi_{0}(G)$ in (38), then there exists $x\in\Omega$ such that $f(x)=0$ and $(G_{x})\geq(H)$.
  • (Additivity) For any two disjoint open $G$-invariant subsets $\Omega_{1}$ and $\Omega_{2}$ with $f^{-1}(0)\cap\Omega\subset\Omega_{1}\cup\Omega_{2}$, one has
    $G\text{\rm-deg}(f,\Omega)=G\text{\rm-deg}(f,\Omega_{1})+G\text{\rm-deg}(f,\Omega_{2}).$
  • (Homotopy) For any $\Omega$-admissible $G$-homotopy, $h:[0,1]\times V\to V$, one has
    $G\text{\rm-deg}(h_{t},\Omega)=\mathrm{constant}.$
  • (Normalization) For any open bounded neighborhood of the origin in an isometric Banach $G$-representation $\mathscr{E}$ with the identity operator $\operatorname{Id}:\mathscr{E}\rightarrow\mathscr{E}$, one has
    $G\text{\rm-deg}(\operatorname{Id},\Omega)=(G).$
[…]
  • (Multiplicativity) For any $(f_{1},\Omega_{1}),(f_{2},\Omega_{2})\in\mathcal{M}^{G}$,
    $G\text{\rm-deg}(f_{1}\times f_{2},\Omega_{1}\times\Omega_{2})=G\text{\rm-deg}(f_{1},\Omega_{1})\cdot G\text{\rm-deg}(f_{2},\Omega_{2}),$
    where the multiplication $\cdot$’ is taken in the Burnside ring $A(G)$.
  • (Recurrence Formula) For an admissible $G$-pair $(f,\Omega)\in\mathscr{M}^{G}(\mathscr{E})$, the $G$-degree (38) can be computed using the following Recurrence Formula:
    $n_{H}=\frac{\deg(f^{H},\Omega^{H})-\sum_{(K)>(H)}{n_{K}\,n(H,K)\,\left|W(K)\right|}}{\left|W(H)\right|},$ (39)
    where $\left|X\right|$ stands for the number of elements in the set $X$ and $\deg(f^{H},\Omega^{H})$ is the Brouwer degree of the map $f^{H}:=f|_{\mathscr{E}^{H}}$ on the set $\Omega^{H}\subset\mathscr{E}^{H}$.
Computation of the $G$-equivariant Leray-Schauder degree. Let’s denote by $\{\mathcal{U}_{i}\}_{i\in\mathbb{N}}$ the set of all irreducible $G$-representations and define the $i$-th basic degree as follows:
$\deg_{\mathcal{U}_{i}}:=G\text{\rm-deg}(-\operatorname{Id},B(\mathcal{U}_{i})).$
Given any isometric Banach $G$-representation with a $G$-isotypic decomposition
$\mathscr{E}=\bigoplus_{i\in\mathbb{N}}\mathscr{E}_{i},$
and any bounded $G$-equivariant linear isomorphism $T:\mathscr{E}\to\mathscr{E}$, the Multiplicativity and Homotopy properties of the $G$-equivariant Leray-Schauder degree, together with Schur’s Lemma implies
$G\text{\rm-deg}(\operatorname{Id}-T,B(\mathscr{E}))=\prod_{i\in\mathbb{N}}G\text{\rm-deg}(T_{i},B(\mathscr{E}_{i}))=\prod_{i\in\mathbb{N}}\prod_{\mu\in\sigma_{-}(T)}\left(\deg_{\mathcal{U}_{i}}\right)^{m_{i}(\mu)}$
where $T_{i}=\operatorname{Id}-T|_{\mathscr{E}_{i}}:\mathscr{E}_{i}\rightarrow\mathscr{E}_{i}$ and $\sigma_{-}(T)$ denotes the real negative spectrum of $T$.
Notice that each of the basic degrees:
$\deg_{\mathcal{U}_{i}}=\sum_{(H)\in\Phi_{0}(G)}n_{H}(H),$
can be practically computed, using the recurrence formula (39), as follows:
$n_{H}=\frac{(-1)^{\dim\mathcal{U}_{i}^{H}}-\sum_{H<K}{n_{K}\,n(H,K)\,\left|W(K)\right|}}{\left|W(H)\right|}.$
The following fact is well-known (see for example [6]).
Lemma A.1.
For any irreducible $G$-representation $\mathcal{U}$, the basic degree $\deg_{\mathcal{U}}\in A(G)$ is an involutive element of the Burnside ring, i.e.
$(\deg_{\mathcal{U}})^{2}=\deg_{\mathcal{U}}\cdot\deg_{\mathcal{U}}=(G).$